### v3

# Evaluating and Utilizing Surrogate Outcomes in Covariate-Adjusted Response-Adaptive Designs

Abstract
Surrogate outcomes have long been studied as substitutes for long-term primary outcomes. However, current surrogate evaluation methods do not directly account for their benefits in updating treatment randomization probabilities in adaptive experiments that aim to learn and respond to treatment effect heterogeneity. In this context, surrogate outcomes can expedite updates to randomization probabilities and thus improve expected outcomes of newly-enrolled participants by enabling earlier detection of heterogeneous treatment effects. We introduce a novel approach for evaluating candidate surrogate outcomes that quantifies both of these benefits in sequential adaptive experiments. We also propose a new Covariate-Adjusted Response-Adaptive design that uses an Online-Superlearner to evaluate and adaptively select surrogate outcomes for updating treatment randomization probabilities during the trial. We further introduce a Targeted Maximum Likelihood Estimation method that addresses dependence in adaptively collected data and achieves asymptotic normality without parametric assumptions. Our design and estimation methods show robust performance in simulations, including those using real trial data. Overall, this framework not only provides a comprehensive way to quantify benefits and select among candidate surrogate outcomes, but also offers a general tool for evaluating various adaptive designs with inferences, providing insights into opportunities and costs of alternative designs that could have been implemented.
Keywords: Adaptive Experimental Design, Causal Inference, Targeted Maximum Likelihood Estimation, Surrogacy
1 Introduction
1.1 Motivation
Surrogate outcomes are often considered as a substitute for the primary outcome when the primary outcome requires a long follow-up time or is expensive to measure. In clinical trials, using surrogate outcomes can result in earlier detection of treatment effects, accelerated drug approvals, and reduced costs (Buyse et al.,, 2000). For example, vaccine-induced immune responses can be surrogate endpoints for protection from HIV infection (Gilbert and Hudgens,, 2008). Surrogate outcomes are also valuable in non-health sectors. For example, in digital experiments conducted by private sector companies, experimenters seek to identify useful surrogates to replace primary outcomes that take time to measure, thereby accelerating new product launches (Duan et al.,, 2021; Richardson et al.,, 2023).
While surrogate outcomes have long been studied for their potential to substitute for long-term primary outcomes, most surrogate evaluation methods do not account for the potential benefits of surrogate outcomes in adaptive randomized trials that aim to learn and respond to treatment effect heterogeneity. In contrast to traditional randomized controlled trials with fixed treatment assignment probabilities, adaptive randomization designs modify the allocation of participants to treatment groups based on accumulated data. This approach allows adaptive randomization to maximize statistical precision or increase the likelihood that participants receive more effective treatments. Among various adaptive designs, Response-Adaptive Randomization (RAR) and Covariate-Adjusted Response-Adaptive (CARA) designs are particularly useful for optimizing treatment assignments (Rosenberger and Lachin,, 2015; Robertson et al.,, 2023). RAR adapts treatment allocation based on previous participants’ outcomes, favoring treatments showing greater promise. CARA designs further incorporates baseline covariates to adjust randomization probabilities in response to heterogeneous treatment effects (Rosenberger et al.,, 2001; Rosenberger and Sverdlov,, 2008; Atkinson et al.,, 2011). This approach is useful when treatment effects and optimal treatments vary among different subgroups without a single treatment that is uniformly optimal for all participants because it both increases the probability that trial participants receive the treatment option most effective for them, and increases precision in estimates of the effect of personalized treatment rules that respond to this heterogeneity.
While CARA adaptive designs offer compelling advantages, they face limitations with outcomes that require long follow-up times. Such delays in observable responses hinder timely adaptation of randomization probabilities. In this context, surrogate outcomes, which are measured earlier than the primary outcome, have the potential to enhance the benefit of adaptive randomization for participants within the experiment directly (i.e. improve the expected outcome of newly-enrolled participants) by allowing for more rapid adaptation of randomization probabilities based on earlier detection of heterogeneous treatment effects. Often, however, there are multiple candidate surrogates, raising the question of which, if any, surrogate is most useful to replace the primary outcome to facilitate the CARA procedure.
Many statistical methods to evaluate surrogate outcomes in traditional clinical trials have not yet been optimized for this sequential decision-making problem posed by CARA experiments. Specifically, in these designs, the timing of the availability of the surrogate outcome plays a critical role, including potential trade-offs between the immediacy of early surrogates and their sensitivity and accuracy in indicating optimal treatments. For example, early proxies of the primary outcome can serve as candidate surrogate outcomes. One possible trade-off arises when these early proxies are partially effective (i.e., imperfect surrogates) in identifying the best personalized treatments. Their use as a surrogate allows for an earlier adaptation of randomization probabilities towards more effective treatments for many participants, compared to waiting for the final outcome. However, using these imperfect early proxies might also skew randomization probabilities towards sub-optimal treatments for some participants. Conversely, waiting for final outcomes (or later and potentially better surrogates) before adapting randomization probabilities risks delaying benefits for early participants and skewing randomization probabilities towards suboptimal treatments for early participants due to insufficient data. The sensitivity of a surrogate to treatment effect heterogeneity is a crucial driver of these trade-offs. A surrogate that is better able to accurately detect heterogeneous treatment effect will lead to a higher probability of participants receiving more effective treatments, compared to one that exhibits less sensitivity.
In short, adaptation of randomization probabilities based on a surrogate, as opposed to the final outcome, has the potential to either help or harm participants in a study. The optimal decision of whether or not to adapt treatment randomization based on a surrogate or to wait for the final outcome, and if using a surrogate, which surrogate to use, is determined by complex trade-offs between the timeliness of the surrogate (i.e., the ability of a candidate surrogate to provide more information earlier about which treatments appear to work best for which patients) and the quality of information a surrogate provides (i.e., the ability of a surrogate to accurately and sensitively detect which treatments truly work best for which patients). Navigating these trade-offs requires both novel approaches for quantifying the benefit of candidate surrogates, and novel designs and estimators that make possible online evaluation of these benefits and dynamic surrogate selection in response during the course of a study.
To address these challenges, we propose a novel approach to surrogate outcome evaluation that accounts for sequential decision-making in CARA experiments. A key contribution is the specification of a target estimand defined by a hypothetical (or counterfactual) adaptive design in which each participant’s treatment assignment is guided by surrogate-derived evidence on optimal individualized treatments from previous enrollees. To estimate this target estimand, we propose a Targeted Maximum Likelihood Estimator (TMLE) (van der Laan et al.,, 2011; van der Laan and Rose,, 2018; van der Laan,, 2008), addressing the technical challenge of making statistical inference on our target parameter based on non-i.i.d data collected from adaptive experiments. We obtain a consistent and asymptotically normal TMLE estimator by leveraging the martingale central limit theorem and maximal inequalities for martingale processes, without relying on parametric assumptions. Our TMLE framework (including our analysis of its asymptotic normality) is generally applicable to help answer other causal questions that would be of interest in regular randomized controlled trials. Specifically, we demonstrate how TMLE utilizes the data from the completed adaptive experiment to estimate the average treatment effect and the expected reward under candidate optimal treatment rules (regimes) (Murphy,, 2003; Luedtke and van der Laan, 2016a, ) for maximizing surrogate outcomes and the final outcome.
More broadly, our framework provides a systematic way to evaluate designs and select from a variety of candidate adaptive designs distinguished by various characteristics, not limited to selection between candidate surrogate outcomes. Although experimenters can choose only one adaptive design at a time, our framework offers a counterfactual reasoning approach to evaluate opportunities and costs of alternative designs, enabling informed selection among multiple candidate designs for sequential decision-making processes in real time.
1.2 Related Work
Surrogate outcomes. Numerous statistical frameworks have been proposed for evaluation of surrogate outcomes, including Prentice’s criteria (Prentice,, 1989), proportion of treatment effect explained (PTE) (Freedman et al.,, 1992; Lin et al.,, 1997), direct and indirect causal effects frameworks (Robins and Greenland,, 1992; Joffe and Greene,, 2009), principal stratification frameworks (Frangakis and Rubin,, 2002; Gilbert and Hudgens,, 2008) and meta-analysis frameworks (Daniels and Hughes,, 1997; Buyse et al.,, 2000). Reviews are in VanderWeele, (2013), Weir and Taylor, (2022), and Elliott, (2023). However, existing methods primarily focus on surrogates’ ability to reflect treatment effects on primary outcomes, without directly quantifying their value in adaptive trial designs, particularly regarding speed and accuracy in identifying heterogeneous treatment effects.
Adaptive Experimental Designs. Adaptive designs are another active research area. In clinical trials, adaptive design methods have become very popular due to their flexibility and efficiency (Chow and Chang,, 2008). These designs enable real-time modifications to key aspects like eligibility criteria, randomization probabilities, and other design features based on sequentially accrued data. In particular, Response-Adaptive Randomization (RAR) adjusts treatment assignment probabilities based on the outcomes and treatments of previously enrolled subjects, aiming to optimize objectives ranging from efficient estimation of treatment effects to enhancing participant welfare by minimizing regret (i.e., proportion of participants assigned to a suboptimal treatment) (Hu and Rosenberger,, 2006; Robertson et al.,, 2023).Covariate-Adjusted Response-Adaptive (CARA) designs further refine this approach by incorporating baseline covariates of subjects into the randomization scheme, allowing for more personalized and effective adaptations (Rosenberger et al.,, 2001; Rosenberger and Sverdlov,, 2008; Atkinson et al.,, 2011). Tracing its origins to the seminal work of Thompson, (1933); Robbins, (1952), the RAR and CARA experimental designs are closely linked to a substantial body of research on Multi-Arm Bandits (MAB) and their derivative, Contextual Bandits (CB) (see Bubeck et al., (2012); Lattimore and Szepesvári, (2020)). A significant portion of this research is dedicated to exploring the efficiency of various algorithms to minimize regret (see the survey in Simchi-Levi and Wang, (2023)).
Using surrogate outcomes to facilitate RAR / CARA designs. The use of surrogate outcomes is also discussed in some literature on RAR designs, as they can be quickly observed when the primary outcome is delayed (Robertson et al.,, 2023). Tamura et al., (1994) uses a short-term response to conduct adaptive randomization to treat outpatients with depressive disorder. Huang et al., (2009) uses a Bayes mixture distribution to model the relationship between surrogates and primary outcomes to facilitate the RAR procedure. McDonald et al., (2023) utilizes a pre-trained Bayesian reward model to run bandits without delay. Yang et al., (2023) considers using an external dataset to construct a surrogate outcome to impute the missing primary outcome to run contextual bandit algorithms. While these approaches allow for earlier deployment of CARA procedure before observing primary outcome data, either relying on prior knowledge, external dataset or generalization assumptions about surrogacy, our framework equips researchers with a tool to dynamically re-evaluate multiple candidate surrogates in their comprehensive utility in CARA design and refine their surrogate choice within the ongoing sequential decision-making environment.
Inference after adaptive experiments. Our estimation method is related to the problem of how to make statistical inference based on data collected from adaptive experiments, which is an important research question considered in the adaptive design literature. For example, Zhang et al., (2007) and Zhu, (2015) established asymptotic results for making statistical inferences from data generated by CARA designs under some parametric models. van der Laan, (2008) developed Targeted Maximum Likelihood Estimation (TMLE) (van der Laan and Rubin,, 2006; van der Laan et al.,, 2011; van der Laan and Rose,, 2018), specifically tailored for adaptive design data, for estimating causal estimand nonparametrically, without relying on specific (e.g. parametric) model assumptions. Following that, Chambaz and van der Laan, (2014) proposed a CARA design targeting an optimal design for estimating the risk difference and log-relative risk, and provides statistical inference robust to model mis-specification with TMLE methodology. Chambaz et al., (2017) established a CARA design targeted for estimating the mean outcome of the optimal treatment rule, and proved a central limit theorem which enables the construction of confidence intervals without relying on parametric model assumptions. Recently, several works used a stabilized doubly robust estimator, following an approach originally proposed by (Luedtke and van der Laan, 2016a, ) to provide inference for the mean outcome under a possibly non-unique optimal treatment strategy, to make valid inference to evaluate treatment effects or the mean outcome under a target treatment rule in non-contextual setting (Hadad et al.,, 2021) and contextual setting (Bibaut et al.,, 2021; Zhan et al.,, 2021).
While existing methodologies focus on inferring the counterfactual mean outcome under a single treatment rule, either predefined or learned from data, we propose a distinct target estimand that captures the timeliness of surrogate outcomes in the sequential decision-making process of CARA procedures. Specifically, our target estimand represents the counterfactual mean outcome across a sequence of stochastic treatment rules optimized to maximize surrogate outcomes, conditioned on finite summaries of historical data before each treatment decision. This approach is related to causal inference for time-series data (van der Laan and Malenica,, 2018; Malenica et al.,, 2021, 2024), where the estimand is defined as an average over a series of time-specific treatment rules based on historical data summaries.
The remainder of this manuscript is organized as follows. Section 2 introduces a general framework that includes data structures, notation, and the statistical estimation setup used in CARA experiments, which lays the groundwork for more detailed discussions on our proposed surrogate evaluation methods and adaptive design mechanisms in Section 3. Section 4 presents an asymptotic analysis of our TMLE estimator. Simulation results are presented in Section 5, followed by an extended simulation study based on real-world trial data in Section 6. Finally, we conclude with discussion and final remarks in Section 7.
2 Overview of Our Design and Estimation Framework
2.1 Overview
In this section, we outline the fundamental data structure, notation, and statistical estimation framework for a sequential Covariate-Adjusted Response-Adaptive (CARA) experiment, in which participants are enrolled, treated, and followed over time and may produce multiple outcomes. While this section provides a brief overview of how surrogates are incorporated and evaluated in these and our proposed designs, we relegate more comprehensive details to Section 3, since the framework to be presented below is generalizable to a wide range of adaptive designs, offering a foundation that applies regardless of the specific mechanism used to update treatment randomization probabilities.
Suppose we have binary treatment $A$ from the set $\mathcal{A}=\{0,1\}$. We use $W\in\mathcal{W}$ to denote the baseline covariates of subjects. Let $Y_{K}$ denote the final outcome we are interested in, where $K$ denotes the follow-up time required to observe $Y_{K}$ after assigning a treatment to each subject. We assume (without loss of generality) that $Y_{K}$ falls within the range of 00 to $1$ and that higher values indicate better outcomes. We also consider $K-1$ shorter-term surrogate outcomes, denoted by $Y_{k}$, each requiring a follow-up time of $k=1,\cdots,K-1$ time points.
Let $T$ represent the total duration of the adaptive experiment. At each time point $t$, the experimenter reviews historical data related to a surrogate outcome $Y_{k}$, estimates the Conditional Average Treatment Effect (CATE) for $Y_{k}$ (represented as $E[Y_{k}|A=1,W]-E[Y_{k}|A=0,W]$ under previously stated identification assumptions) and then generates a treatment randomization function $\tilde{g}_{k,t}:\mathcal{A}\times\mathcal{W}\to[0,1]$ that maps treatment levels and covariate values to a treatment randomization probability based on the estimated CATE for $Y_{k}$ and the covariates of newly-enrolled participants at time $t$ in the sense that they have higher chance to be assigned the better personalized treatment. Details on how we estimate CATE functions and adapt treatment randomization probabilities based on these estimate in the surrogate-based design will be introduced in Section 3.1.
In order to address unknown trade-offs implied by different candidate surrogates in adaptive designs, we will later introduce a novel adaptive design in Section 3.2 that continually evaluates each surrogate’s utility (via a target estimand with statistical estimation and inference) without relying on prior knowledge. This target estimand quantifies the expectation of final outcome $Y_{K}$ assuming every participant were treated in a hypothetical (or counterfactual) adaptive design that was responsive to optimizing a surrogate outcome at the time of their enrollment. We use Targeted Maximum Likelihood Estimation (TMLE, Section 3.3) to estimate the target estimands and identify the surrogate with the largest lower confidence bound of the TMLE estimate. This surrogate is then utilized to adjust treatment randomization probabilities for newly enrolled subjects (Section 3.4). This approach enhances the design’s adaptivity and effectiveness by selecting surrogates that are more likely to yield benefits for subjects within the experiment.
In the remainder of this section, we first introduce more comprehensive notations for the statistical setup, including observed data and its underlying causal model. Then we define a general target estimand under a counterfactual CARA design. These foundations will later support the development of target estimands for evaluating surrogates and an in-depth analysis of surrogate-based adaptive designs in Section 3.
2.2 Statistical Setup
We use $t=1,\cdots,T$ to index the time that a participant is initially enrolled in the study and initially randomized, and $k$ to index the on-study follow-up time of each participant. Let $E(t)$ denote the number of subjects who enroll at each time point $t$, and define $N(t)=\sum_{t^{\prime}=1}^{t}E(t^{\prime})$ as the cumulative number of participants enrolled up to time $t$. For each participant $i$, we observe $O_{i}=(t_{i},W_{i},A_{i},Y_{i,1},\cdots,Y_{i,K})$ where $t_{i}$ is entry time; $W_{i}$ and $A_{i}$ are the baseline covariates and treatment, observed at time $t_{i}$; at time $t_{i}+k$ ($k=1,\cdots,K-1$), we observe $K-1$ surrogate outcomes $Y_{i,1},...,Y_{i,K-1}$ corresponding to follow-up times from $k=1$ to $K-1$. At time $t_{i}+K$, we observe $Y_{i,K}$, the final outcome of interest of the $i$-th participant. In order to emphasize the chronological time that each node is observed, we define $\tilde{W}_{i}(t_{i})\equiv W_{i},\tilde{A}_{i}(t_{i})\equiv A_{i}$, and $\tilde{Y}_{i}(t_{i}+k)\equiv Y_{i,k}$ for $k=1,\cdots,K$. We use the two sets of notations interchangeably. We then denote the historical data from time $1$ to $t-1$ by $\bar{O}(t-1)=\{{\tilde{W}_{i}(t^{\prime}):i\in[N(t-1)],t^{\prime}\leq t-1}\}\bigcup\{\tilde{A}_{i}(t^{\prime}):i\in[N(t-1)],t^{\prime}\leq t-1\}\bigcup\{\tilde{Y}_{i}(t^{\prime}):i\in[N(t-1)],t^{\prime}\leq t-1\}$. As time moves from $t-1$ to $t$, we will observe more nodes from two groups of subjects: $R_{1}(t)=\{i:t-K\leq t_{i}\leq t-1\}$ and $R_{2}(t)=\{i:t_{i}=t\}$. $R_{1}(t)$ are all previously enrolled participants who haven’t reached their endpoint before time $t$; $R_{2}(t)$ are all participants who enter the experiment beginning at time $t$. At each time $t$, we first observe outcomes $Y_{i,t-t_{i}}$ for all $i\in R_{1}(t)$, subsequently observe covariates $W_{j}$, and then assign treatment $A_{j}$ for each newly-enrolled participants $j\in R_{2}(t)$.
We use ¯$\overline{\mathcal{H}}(t)=\bar{O}(t-1)\bigcup\{\tilde{Y}_{i}(t),i\in R_{1}(t)\}$ to denote the historical data before the observation of $\{(W_{j},A_{j}),j\in R_{2}(t)\}$. We define $\tilde{g}_{t}:\mathcal{A}\times\mathcal{W}\to[0,1]$ as a time $t$-specific function that presents the conditional distributions of $A$ given $W$. Let $\mathcal{G}_{t}$ denote a set of conditional distributions of $A$ given $W$ that includes $\tilde{g}_{t}$. We note that $\tilde{g}_{t}$ is itself an output of a mapping from ¯$\overline{\mathcal{H}}(t)$ to $\mathcal{G}_{t}$. In an adaptive experiment, $\tilde{g}_{t}$ is defined based on historical data ¯$\overline{\mathcal{H}}(t)$, and subsequently used to assign treatment randomization probabilities to every new unit $j\in R_{2}(t)$ based on their baseline covariates $W_{j}$.
To formalize this for individual participants within the experiment, we define $\mathcal{H}(i):=(\overline{\mathcal{H}}(t_{i}),W_{i})$ to denote the unit $i$-specific historical data before treatment $A_{i}$. We focus on an $i$-specific stochastic rule that depends on $\mathcal{H}(i)$ through a fixed dimensional summary measure $C_{i}:=f_{C}(\mathcal{H}(i))\in\mathcal{C}$, where $f_{C}$ is a fixed function that maps the historical data $\mathcal{H}(i)$ into a fixed-dimensional real-valued space. In practice, we require that $C_{i}$ always includes the baseline covariates $W_{i}$. Given the summary measure $C_{i}$ before $A_{i}$, we define $g_{i}(\cdot|C_{i}):\mathcal{A}\to[0,1]$ as the function of assigning treatment randomization probability for any $a\in\mathcal{A}=\{0,1\}$.
Let $q_{W}$ denote the marginal density of $W$, and $q_{Y_{k}}$ denote the conditional density of outcome $Y_{k}$ given $W,A,Y_{1},\cdots,Y_{k-1}$ ($k=1,\cdots,K$). We can represent the likelihood of data $\bar{O}(t)$ by: $p\left(\bar{o}(t)\right)=\prod_{i=1}^{N(t)}q_{W}(w_{i})g_{i}(a_{i}|C_{i})\times\prod_{k=1}^{\min(K,t-t_{i})}q_{Y_{k}}(y_{i,k}|w_{i},a_{i},y_{i,1},\cdots,y_{i,k-1})$.
We define $\bar{Q}_{k}(A,W)=E(Y_{k}|A,W)$ to be the conditional mean for $k=1,\cdots,K$, and define $\theta=(g,q_{w},\bar{Q}_{1},\cdots,\bar{Q}_{K})\in\mathcal{G}\times\mathcal{Q}_{W}\times\bar{\mathcal{Q}}_{1}\times\cdots\times\bar{\mathcal{Q}}_{K}$. $\bar{Q}_{1},\cdots,\bar{Q}_{K}$ are common functions across all units. Conditional on the realized summary $C_{i}$ of every unit $i$, the statistical model of distributions of $O_{i}$ is $\mathcal{M}(C_{i})=\{P_{\theta,C_{i}}:\theta\}$.
Let $\bar{O}_{i}(t)$ denote the cumulative data $O_{i}$ observed up to time $t$ for the $i$-th participant. We can rewrite the conditional likelihood of $\bar{O}_{i}(t)$ up to time $t$ given the observed summary measure $C_{i}$ as: $p_{\theta,C_{i}}(o_{i})=q_{W}(w_{i})g_{i}(a_{i}|C_{i})\prod_{k=1}^{t-t_{i}}q_{Y_{k}}(y_{i,k}|w_{i},a_{i},y_{i,1},\cdots,y_{i,k-1})$.
We specify a Structural Causal Model (SCM) (Pearl et al.,, 2000) for variables of new unit $i$ enrolled at time $t_{i}$ by $W_{i}=f_{W}(U_{W_{i}});\,\mathcal{H}(i)=(\overline{\mathcal{H}}(t),W_{i});\,C_{i}=f_{C}(\mathcal{H}(i));\,A_{i}=f_{A}(C_{i},U_{A_{i}});\,Y_{i,k}=f_{Y_{k}}(W_{i},A_{i},Y_{i,1},\cdots,Y_{i,k-1},U_{Y_{i,k}}),\,k=1,\cdots,K$. Here, we denote unmeasured and independent exogenous variables by $U_{i}=(U_{W,i},U_{A,i},U_{Y_{i,1}},\cdots,U_{Y_{i,K}})$, which follows the same distribution across all units. We use $g_{0,i}$ to denote the actual stochastic rule for assigning treatments for the $i$-th participant, and $\mathbf{g}_{0,1:n_{t}}=(g_{0,1},\cdots,g_{0,n_{t}})$ denote the actual adaptive design used for treatment assignment and data collection.
2.3 Target Estimand under Counterfactual CARA Design
We first propose a general target estimand of interest under a modified SCM under any counterfactual adaptive design. We start by defining a vector of summary measures $\mathcal{\bar{C}}^{*}(n):=(C_{1}^{*},\cdots,C_{n}^{*})$, which are generated sequentially in a counterfactual adaptive design denoted by $\mathbf{g}_{*,1:n}(\mathcal{\bar{C}}^{*}(n)):=(g_{*,i}(\cdot|C_{i}^{*}))_{i=1,\cdots,n}$, where each $g_{*,i}(\cdot|C_{i}^{*})$ is a stochastic treatment rule $\mathcal{A}\to[0,1]$, depending on $C_{i}^{*}$. The generations of $\mathcal{\bar{C}}^{*}(n)$ and $\mathbf{g}_{*,1:n}(\mathcal{\bar{C}}^{*}(n))$ are illustrated in the following modified SCM: $W_{i}=f_{W}(U_{W}),\mathcal{H}^{*}(i)=(\overline{\mathcal{H}}^{*}(t_{i}),W_{i}),C_{i}^{*}=f_{C}(\mathcal{H}^{*}(i)),A_{i}^{*}\sim g_{*,i}(\cdot|C_{i}^{*}),Y_{i,k}^{*}=f_{Y_{k}}(W_{i},A_{i}^{*},Y_{i,1}^{*},\cdots,Y_{i,k-1}^{*},U_{Y_{k}}),k=1,\cdots,K$, in which $f_{A}$ in the original SCM are replaced by stochastic rules $g_{*,i}(\cdot|C_{i}^{*})$. We note that each element $C_{i}^{*}$ influences the subsequent treatment rule $g_{*,i}(\cdot|C_{i}^{*})$, thus impacting the subsequent $A_{i}^{*}$, $Y_{i,k}^{*}$, and $C_{i+1}^{*}$. This model underscores the iterative and cumulative impact had we intervened with a series of counterfactual stochastic treatment rules.
An ideal target estimand of interest is the counterfactual expectation of $Y_{K}^{*}$ under the counterfactual adaptive design $\mathbf{g}_{*,1:n}(\mathcal{\bar{C}}^{*}(n))$ generated by that modified SCM, marginalizing over all possible realizations of $C_{i}^{*}$. While it is theoretically feasible to identify this target estimand by using the $g$-computation formula (Robins,, 1986, 1987), it results in a complex statistical estimand, for which any estimation procedure would be computationally intensive in the nested data structure where any small change in early nodes could have a cumulative impact on follow-up nodes.
Alternatively, we can modify our target estimand by following the causal inference framework for single time-series data, as introduced in van der Laan and Malenica, (2018) and Malenica et al., (2021). We set up another SCM that replaces the original treatment generating process with $A_{i}^{*}\sim g_{*,i}(\cdot|C_{i})$, followed by $Y_{i,k}^{*}=f_{Y_{k}}(W_{i},A_{i}^{*},Y_{i,1}^{*},\cdots,Y_{i,k-1}^{*},U_{Y_{k}})$ ($k=1,\cdots,K$), where our $g_{*,i}$ is defined by the observed summary measure $C_{i}$ for every unit $i$. Under this simplified SCM with less modification, we propose the following data-adaptive parameter conditional on $\mathcal{\bar{C}}(n):=(C_{1},\cdots,C_{n})$:
$\Psi_{\mathbf{g}_{*,1:n}\left(\mathcal{\bar{C}}(n)\right)}=\frac{1}{n}\sum_{i=1}^{n}E_{g_{*,i}\left(\cdot|C_{i}\right)}\left[Y_{i,K}^{*}\right],$
which targets the counterfactual final outcome of all units had they been treated, at their entry time, by different stochastic treatment rules given the summary measures observed from the experiment.
While comparing these two estimands is itself an interesting research topic, for the rest of this paper we focus on evaluating counterfactual adaptive designs using $\Psi_{\mathbf{g}_{*,1:n}\left(\mathcal{\bar{C}}(n)\right)}$ as the target estimand, since it allows for robust estimation under sequential adaptive designs with reduced computational burden. Detailed discussions of their similarities and differences are provided in Appendix A. Our simulation results also reveal that the numerical differences between the estimands are minimal in our context.
3 Evaluating and Utilizing Surrogates in CARA design
In this section, we first introduce how we design a CARA procedure to adapt treatment randomization probabilities based on surrogates. Building upon the framework in Section 2, we define our target estimand that evaluates each candidate surrogate within a counterfactual adaptive design that utilizes the surrogate to adapt treatment randomization probabilities. We will also demonstrate how to use the TMLE to estimate these target estimands to evaluate surrogates. Finally we will introduce a new CARA design that incorporates an Online Superlearner to identify and utilize the most useful surrogate.
3.1 CARA Designs Using Surrogates Outcomes
We focus on a class of CARA designs that utilize a candidate surrogate as a response to adapt treatment randomization probabilities, following an extension of the strategy proposed by Chambaz et al., (2017). Our strategy balances exploration and exploitation based on the CATE functions learned from the data. Specficially, $g^{*}_{k,i}(a|C_{i})$ depends on the confidence interval of the estimated CATE of $Y_{k}$ for participant $i$. The more deviated the lower bound of the confidence interval is from zero, the more likely it is to assign unit $i$ to an optimal individual treatment. A minimum probability (e.g., 10%) is designated for the seemingly inferior treatment learned from data of earlier participants, while a maximum (e.g., 90%) is set for the seemingly superior one. If the confidence interval for the estimated CATE excludes zero, the superior treatment is assigned with the maximum allowable probability. If no outcome $Y_{k}$ has been observed prior to the enrollment of a subject, we assign equal randomization probabilities of 0.5 for both treatments. We refer to Appendix D for details on the CARA procedure to generate of treatment randomization probabilities.
3.2 Target Estimand for Evaluating Surrogates in CARA Designs
As a specification of the general target estimand under a hypothetical adaptive design (as introduced in Section 2.3), we consider $K$ candidate adaptive designs $\mathbf{g}^{*}_{k,1:n_{t}}$ ($k=1,\cdots,K$) for all units enrolled from time $1$ to $t-K$ with complete follow-up, using $n_{t}=N(t-K)$ to denote the total number of units till time $t-K$. In the $k$-th design $\mathbf{g}^{*}_{k,1:n_{t}}$, the stochastic rule $g^{*}_{k,i}$ is generated by the estimated conditional average treatment effect on the surrogate outcome $Y_{k}$ using nodes $W$, $A$ and $Y_{k}$ observed from historical data $\mathcal{H}(i)$. We define the following target estimand for each surrogate $Y_{k}$:
$\Psi_{t,k}:=\Psi_{t,\mathbf{g}^{*}_{k,1:n_{t}}}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}E_{g^{*}_{k,i}}\left[Y_{i,K}^{*}\right].$
We refer to this parameter of “the utility of a candidate surrogate” in the CARA design. Our target estimand corresponds to the expected final outcome $Y_{K}$ of all $n_{t}$ units had they been treated, at their entry time, by a stochastic rule tailored to optimize the mean outcome of an early surrogate $Y_{k}$. We will use this target estimand to evaluate the utility of different surrogate outcomes in the CARA design, and utilize this to choose the surrogate to be used for updating the treatment allocation strategy used for the next time point in the ongoing adaptive design.
Let $Q_{0}:=(\bar{Q}_{0,K},Q_{0,W})\in\bar{\mathcal{Q}}_{K}\times\mathcal{Q}_{W}$, where $Q_{0,W}$ is the true distribution of $W$, and $\bar{Q}_{0,K}$ is the true conditional mean function of the final outcome $Y_{K}$ given $A$ and $W$. We can identify our causal estimand (i.e., the utility of a candidate surrogate) by the observed data under the following assumptions. Since the investigators are designing the adaptive experiments, they can carefully specify their designs to ensure that the causal identification assumptions hold.
Assumption 1 (Sequential Exchangeability).
For any $k\in[K]$ and any subject $i\in\{1,\cdots,n_{t}\}$, $Y_{i,k}(a)\perp A_{i}\mathrel{|}C_{i}$.
Assumption 2 (Positivity).
For any subject $i$, $g_{0,i}(a_{i}|C_{i})>0$ for any $a_{i}\in\mathcal{A}$.
Theorem 1 (Identification).
With assumptions 1 and 2, one can identify $\Psi_{t,k}$ under $Q_{0}$ from the observed data by
$\Psi_{t,k}(Q_{0})=\Psi_{t,\mathbf{g}^{*}_{k,1:n_{t}}}(Q_{0})=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}E_{g^{*}_{k,i},Q_{0}}\left[Y_{i,K}^{*}\right]=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\sum_{a=0}^{1}\bar{Q}_{0,K}(a,w_{i})g^{*}_{k,i}(a|C_{i})\\ .$
3.3 Targeted Maximum Likelihood Estimation
We propose using a Targeted Maximum Likelihood Estimator (TMLE) to estimate $\Psi_{t,k}$. Let $\bar{Q}_{n,K,t}$ be an initial estimate of $\bar{Q}_{0,K}$ based on the historical data $\bar{H}(t)$ observed up till time $t$. One can employ the Super Learner (van der Laan et al.,, 2007) or Highly Adaptive Lasso (van der Laan,, 2015, 2017; Benkeser and van der Laan,, 2016; van der Laan,, 2023) to obtain an initial estimator $\bar{Q}_{n,K,t}$ of the conditional mean function $\bar{Q}_{0,K}$. Then we apply a TMLE procedure to update $\bar{Q}_{n,K,t}$ by fitting the following logistic model:
$\mathop{\mathrm{logit}}\bar{Q}_{n,K,t,\epsilon}\left(A,W\right)=\mathop{\mathrm{logit}}\bar{Q}_{n,K,t}\left(A,W\right)+\epsilon$
with weights $\frac{g^{*}_{k,i}(a_{i}|C_{i})}{g_{0,i}(a_{i}|C_{i})}$ (van der Laan and Rubin,, 2006; van der Laan,, 2008; van der Laan et al.,, 2011; Malenica et al.,, 2024). The estimate of $\epsilon$ is denoted by $\hat{\epsilon}_{t}$, and an updated estimator of the conditional mean function, $\bar{Q}_{n,K,t}^{*}:=\bar{Q}_{n,K,t,\hat{\epsilon}_{t}}$, is obtained, which solves the following estimating equation: $\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\frac{g^{*}_{k,i}(a_{i}|C_{i})}{g_{0,i}(a_{i}|C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,t}^{*}(a_{i},w_{i})\right)=0$. The TMLE estimate for $\Psi_{t,k}$ is thus given by
$\hat{\Psi}_{t,k}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\sum_{a=0}^{1}\bar{Q}_{n,K,t}^{*}(a,w_{i})g^{*}_{k,i}(a|C_{i}).$
We estimate the standard error of $\hat{\Psi}_{t,k}$ by $\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}}$, where
$\hat{\sigma}^{2}_{t,k}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\left[\frac{g^{*}_{k,i}(a_{i}|C_{i})}{g_{0,i}(a_{i}|C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,t}^{*}(a_{i},w_{i})\right)\right]^{2}.$
We also provide the following confidence interval for $\hat{\Psi}_{t,k}$:
$\left[\hat{\Psi}_{t,k}-z_{1-\alpha/2}\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}},\hat{\Psi}_{t,k}+z_{1-\alpha/2}\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}}\right],$
where $z_{1-\alpha/2}$ is the $21-\alpha/$ 2 quantile of standard normal distribution ($\alpha=0.05$). This method is widely applicable for estimating the target estimand $\Psi_{\mathbf{g}_{*,1:n}\left(\mathcal{\bar{C}}(n)\right)}$ under a general counterfactual adaptive design $\mathbf{g}_{*,1:n}\left(\mathcal{\bar{C}}(n)\right)$, as introduced in Section 2.3.
3.4 Online Superlearner that Selects Surrogates in CARA Design
As we progress to a new time point $t$, we have up to $K$ candidate stochastic rules $g^{*}_{1,i},\cdots,g^{*}_{K,i}$ for assigning treatment randomization probabilities for a new participant $i$ enrolled at $t$. For each $k$, the stochastic rule $g^{*}_{k,i}$ is generated based on historical data, specifically tailored to optimize the mean outcome of $Y_{k}$. We next propose the use of an Online Superlearner to select the most useful surrogate at time $t$ (denoted as $Y_{k^{*}_{t}}$), which is used to guide subsequent treatment randomization decisions; importantly, this approach allows surrogate selection to depend on information about a surrogate’s utility accrued during the study, and allows this surrogate selected to evolve over the course of the study. This process begins by estimating $\hat{\Psi}_{t,k}$ for each $k\in[K]$ and obtaining confidence intervals for these estimates. The surrogate $Y_{k^{*}_{t}}$ is then determined by identifying $k^{*}_{t}$, which is defined as:
$k^{*}_{t}=\mathop{\mathrm{argmax}}_{k\in[K]}\hat{\Psi}_{t,k}-z_{1-\alpha/2}\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}},$
where $k^{*}_{t}$ corresponds to the surrogate with the highest lower confidence interval bound among all $\hat{\Psi}_{t,k}$ for $k=1,\cdots,K$. Finally, for all units $i$ enrolled at time $t$, we assign treatment randomization probabilities using $g^{*}_{k_{t}^{*},i}$.
3.5 Statistical Inference at the End of CARA Experiments
When the experiment ends at time $T$, one can follow the same methodology outlined in Section 3.3 to estimate $\Psi_{T,k}$. Furthermore, we note that the non-i.i.d. data collected from our adaptive experiments also allow us to identify and estimate (with rigorous inference) other causal estimands typically studied in randomized controlled trials. These include, but are not limited to, the average treatment effect and the expected reward under a learned treatment rule designed to optimize candidate surrogate outcomes or the final outcome. For these target estimands, a doubly-robust TMLE estimator is feasible, specifically designed to handle the challenges of data dependency while ensuring asymptotic normality. The methodology for constructing TMLEs for these estimands mirrors the approach of making statistical inference for our target estimand $\Psi_{T,k}$ and the other estimands shown in van der Laan, (2008); Chambaz et al., (2017); Bibaut et al., (2021); Malenica et al., (2021, 2024).
For example, to estimate the average effect of treatment on the primary outcome $Y_{K}$ across all participants, defined as $\psi_{n,K,T}^{\text{ATE}}:=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\left[\bar{Q}_{0,K}(1,w_{i})-\bar{Q}_{0,K}(0,w_{i})\right]$, one can start by obtaining an initial estimator $\bar{Q}_{n,K,T}:\mathcal{A}\times\mathcal{W}\to[0,1]$ for the conditional mean function $\bar{Q}_{0,K}$. This is followed by applying a TMLE update to derive $\bar{Q}_{n,K,T}^{*}$, which solves the equation $\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\frac{2a_{i}-1}{g_{0,i}(a_{i}|C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,T}^{*}(a_{i},w_{i})\right)=0.$ The TMLE estimate of ATE is then $\hat{\psi}_{n,K,T}^{\text{ATE}}:=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\left[\bar{Q}_{n,K,T}^{*}(1,w_{i})-\bar{Q}_{n,K,T}^{*}(0,w_{i})\right]$. The standard error is calculated by $\sqrt{\frac{1}{n_{T}}\hat{\sigma}^{2,\text{ATE}}_{n,k,T}}$, where $\hat{\sigma}^{2,\text{ATE}}_{n,k,T}=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\left[\frac{2a_{i}-1}{g_{0,i}(a_{i}\mid C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,T}^{*}(a_{i},w_{i})\right)\right]^{2}.$
Additionally, given an estimate $d^{*}_{n,k,T}:\mathcal{W}\to\mathcal{A}$ of the true optimal treatment rule $d_{0,k}$ for maximizing the expected mean of a surrogate outcome $Y_{k}$, we define the target estimand $\psi_{d^{*}_{n,k,T}}=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\bar{Q}_{0,K}(d^{*}_{n,k,T}(w_{i}),w_{i})$ , which represents the mean final outcome $Y_{K}$ under an estimated optimal treatment rule for a surrogate $Y_{k}$. To estimate $\psi_{d^{*}_{n,k,T}}$, we first compute an initial estimator $\bar{Q}_{n,K,T}$ and update it by using the TMLE procedure to satisfy $\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\frac{d^{*}_{n,k,T}(w_{i})}{g_{0,i}(a_{i}|C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,T}^{*}(a_{i},w_{i})\right)=0.$ With $\bar{Q}_{n,K,T}^{*}$, the TMLE estimate of $\psi_{d^{*}_{n,k,T}}$ is given by: $\hat{\psi}_{d^{*}_{n,k,T}}=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\bar{Q}_{n,K,t}^{*}(d^{*}_{n,k,T}(w_{i}),w_{i}).$ The standard error is estimated by $\sqrt{\frac{1}{n_{T}}\hat{\sigma}^{2}_{d^{*}_{n,k,T}}}$, where $\hat{\sigma}^{2}_{d^{*}_{n,k,T}}=\frac{1}{n_{T}}\sum_{i=1}^{n_{T}}\left[\frac{d^{*}_{n,k,T}(w_{i})}{g_{0,i}(a_{i}|C_{i})}\left(Y_{i,K}-\bar{Q}_{n,K,T}^{*}(a_{i},w_{i})\right)\right]^{2}.$ However, since $d^{*}_{n,K,T}$ is an optimal treatment rule trained to maximize the primary outcome $Y_{K}$ itself, applying regular TMLE estimate of the mean primary outcome under this rule may exhibit upward bias. This bias occurs because the same data is used both to learn the rule and to estimate its performance. We recommend using cross-validated TMLE (CV-TMLE) (Zheng and van der Laan,, 2010; van der Laan and Luedtke,, 2015; Montoya et al.,, 2021), which uses internal sample splitting to avoid finite sample bias.
We note that $\psi_{d^{*}_{n,k,T}}$ can be an interesting supplemental target estimand to evaluate a surrogate, as it reflects the mean primary outcome under the surrogate-based optimal treatment rule. This type of approach was also introduced in Hsu et al., (2015) and Yang et al., (2023). However, similar to most existing surrogacy evaluation methods and in contrast to the surrogate utility parameter we propose here, this measure does not evaluate the timeliness and sensitivity of surrogates in detecting effective treatments to benefit participants more rapidly.
4 Theoretical Analysis of TMLE
In this section, we show consistency and asymptotic normality of the proposed TMLE $\hat{\Psi}_{t,k}$ by decomposing its difference from the truth $\Psi_{t,k}$ into a sum of martingale difference sequence as the first term and a martingale process as the second term. We analyze them by Martingale Central Limit Theorem and equicontinuity conditions.
For any $\bar{Q}_{K}\in\bar{\mathcal{Q}}_{K}$, we define $D_{k}(\bar{Q}_{K})\left(O_{i},C_{i}\right)=\frac{g^{*}_{k,i}\left(A_{i}|C_{i}\right)}{g_{0,i}\left(A_{i}|C_{i}\right)}\left(Y_{i,K}-\bar{Q}_{K}\left(A_{i},W_{i}\right)\right)$, where $C_{i}$ is a summary measure, including baseline covariates $W_{i}$. {restatable}theoremrepThmDecomposition For any $\bar{Q}_{n,K,t}^{*}\in\bar{\mathcal{Q}}_{K}$ and $\bar{Q}_{1,K}\in\bar{\mathcal{Q}}_{K}$, the difference between $\hat{\Psi}_{t,k}$ and $\Psi_{t,k}$ can be decomposed as $\hat{\Psi}_{t,k}-\Psi_{t,k}=M_{1,n_{t}}(\bar{Q}_{1,K})+M_{2,n_{t}}(\bar{Q}_{n,K,t}^{*},\bar{Q}_{1,K}),$ where
$M_{1,n_{t}}\left(\bar{Q}_{1,K}\right)$ $=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\left[D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)-E\left[D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)|\mathcal{H}(i)\right]\right],$
$M_{2,n_{t}}\left(\bar{Q}_{n,K,t}^{*},\bar{Q}_{1,K}\right)$ $=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\Biggl{[}\biggl{[}D_{k}\left(\bar{Q}_{n,K,t}^{*}\right)\left(O_{i},C_{i}\right)-E\left[D_{k}\left(\bar{Q}_{n,K,t}^{*}\right)\left(O_{i},C_{i}\right)|\mathcal{H}(i)\right]\biggr{]}$
$-\biggl{[}D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)-E\left[D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)|\mathcal{H}(i)\right]\biggr{]}\Biggr{]}.$
4.1 Analysis of the First Term
We first analyze the first term $\sqrt{n_{t}}M_{1,n_{t}}(\bar{Q}_{1,K})$, which is a sum of martingale difference sequence. One can apply Martingale central limit theorem to prove its asymptotic normality.
Assumption 3 (Strong positivity).
For every $i$, there exists $\zeta>0$ such that $g_{0,i}>\zeta$ and $g^{*}_{k,i}>\zeta$.
Assumption 4 (Stabilization of the average of conditional variances).
[…]
Theorem 2 (Asymptotic normality of the first term).
[…]
Proof.
The result directly follows from various forms of the martingale central limit theorems (e.g. Theorem 2 in Brown, (1971)). ∎
4.2 Analysis of the Second Term
Now we suppress the notation $t$ by denoting the sample size $n_{t}$ by $N$, the estimated conditional mean of final outcome $\bar{Q}_{n,K,t}^{*}$ by $\bar{Q}_{N,K}^{*}$ and the equicontinuity term $M_{2,n_{t}}$ by $M_{2,N}$. To construct an asymptotically normal estimator, we need the negligibility of the second term $M_{2,N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})$, i.e., $M_{2,N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})=o_{P}(N^{-\frac{1}{2}}).$
[…]
Assumption 5 (Covering Number of $\bar{Q}_{K}$).
$\bar{\mathcal{Q}}_{K}$ is a class of functions with finite sup-norm covering number $N_{\infty}({\epsilon},\bar{\mathcal{Q}}_{K})$.
Assumption 6 (Convergence of $\bar{Q}_{N,K}^{*}$).
Define
$s_{N}\left(\bar{Q}_{K},\bar{Q}_{1,K}\right):=\sqrt{\frac{1}{N}\sum_{i=1}^{N}E\left[\phi\left(\frac{|D_{k}\left(\bar{Q}_{K}\right)(O_{i},C_{i})-D_{k}\left(\bar{Q}_{1,K}\right)(O_{i},C_{i})|}{Z}\right)\bigm{|}C_{i}\right]},$
where $\phi(x):=e^{x}-x-1$, $Z:=\frac{16(1-\zeta)^{2}}{\zeta^{2}}$. Assume that $s_{N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})=o_{P}(1).$
Theorem 3 (Equicontinuity).
Under Assumptions 3, 5, and 6,
$M_{2,N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})=o_{P}(N^{-\frac{1}{2}}).$
The proof depends on general results in Appendix E, which proves equicontinuity of martingale process in a general setup (see Theorem 6 and Corollary 6). The proof of Theorem 3 for adaptive design is in Appendix F.
Recent work of van der Laan, (2023) introduces a class of $m$-th order smoothness functions with bounded sectional variation norm on up to their $m$-th-order derivatives and the higher order spline Highly Adaptive Lasso (HAL) estimator, which will be reviewed in Appendix B. The theoretical results in van der Laan, (2023) imply that if $\bar{Q}_{1,K}$ is in that class of functions and $\bar{Q}_{N,K}$ is estimated by higher order spline Highly Adaptive Lasso (HAL), we can prove the equicontinuity condition without relying on Assumption 6. More details about these results are discussed in Appendix E.
{restatable}
assumptionrepAssmpHigherOrderSpline $\bar{Q}_{1,K}$ and $\bar{Q}_{N,K}^{*}$ are from a class of functions that are represented by an $m$-th order ($m=1,2,\cdots$) smooth function whose $m$-th order derivatives are Càdlàg functions with bounded variation norm (van der Laan,, 2023). {restatable}corollaryrepCoroHighOrderSplineEquicontinuity Suppose Assumptions 3 and 3 hold. Then $M_{2,N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})=o_{P}(N^{-\frac{1}{2}})$.
Proof.
We leave the proof of Theorem 3 to Appendix 3, which depends on the general result of Corollary 6 in Appendix E. ∎
4.3 Asymptotic Normality of TMLE
Theorem 4 (Asymptotic normality of the TMLE).
[…]
The asymptotic normality is directly obtained from the asymptotic normality of the first term $\sqrt{n_{t}}M_{1,n_{t}}(\bar{Q}_{1,K})$ and the equicontinuity condition in the second term $M_{2,n_{t}}\left(\bar{Q}_{n,K,t}^{*},\bar{Q}_{1,K}\right)=o_{P}\left(n_{t}^{-\frac{1}{2}}\right)$. Note that convergence of $\bar{Q}_{N,K}^{*}$ to $\bar{Q}_{0,K}$ is not required. We estimate the standard error $\sigma^{2}_{t,k}$ by $\hat{\sigma}^{2}_{t,k}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\left\{\frac{g^{*}_{k,i}\left(A_{i}\mid C_{i}\right)}{g_{0,i}\left(A_{i}\mid C_{i}\right)}\left[Y_{i,K}-Q_{n,K,t}^{*}\left(A_{i},W_{i}\right)\right]\right\}^{2}$ (which converges in probability to $\sigma^{2}_{t,k}$), and provide a confidence interval as follows:
$\Bigg{[}\hat{\Psi}_{t,k}-z_{1-\alpha/2}\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}},\,\hat{\Psi}_{t,k}+z_{1-\alpha/2}\sqrt{\frac{1}{n_{t}}\hat{\sigma}^{2}_{t,k}}\Bigg{]}.$
5 Simulations
5.1 Simulation Setup
In this section, we present simulation studies aimed at evaluating the performance of our proposed adaptive design and estimation methods. We consider a binary treatment $A\in\{0,1\}$, a one-dimensional baseline covariate $W$, which follows a uniform distribution $U(-4,4)$, four surrogate outcomes $Y_{k}$ measured at follow-up times $k=1,2,3,4$, respectively, and a final outcome $Y_{5}$ measured at follow-up time 5. Each participant’s outcome $Y_{k}$ is generated by its conditional mean $E(Y_{k}|A,W)$ plus a random noise following the normal distribution $N(0,1)$. For detailed information on our design configurations, we refer readers to Appendix C for CATE estimation, Appendix D for CARA mechanism and Appendix G for the estimation of conditional mean functions.
We evaluate the performance of various designs in two scenarios. In Scenario 1, the conditional mean functions of $Y_{k}$ given $A$ and $W$ are $E[Y_{k}|A,W]=(2A-1)\times\left[0.5-(1+\text{exp}(-(3-k)-W))^{-1}\right]$. The CATE function is shown in the left of Figure 1. In this scenario, the optimal treatment rule (based on the final outcome of interest) is more aligned with the optimal treatment rule defined in terms of later versus earlier surrogates. This implies that, despite requiring some time to measure and thus delaying adaptation in randomization probabilities, later surrogates are more useful in guiding treatment assignment probabilities so that participants are more likely to be treated by the correct individualized optimal treatment that maximizes the final outcome.
In Scenario 2, the conditional mean functions of $Y_{k}$ are $E[Y_{k}|A,W]=(2A-1)\times\left[0.5-(1+\text{exp}(-\gamma_{k}W))^{-1}\right]$, where $\gamma_{1}=3$,$\gamma_{2}=2$,$\gamma_{3}=1$,$\gamma_{4}=0.5$ and $\gamma_{5}=0.25$. The CATE functions are shown in the right of Figure 1. In this scenario, all surrogate outcomes have the same optimal treatment rule as that of the final outcome. Notably, early surrogates demonstrate a larger absolute value in the conditional average treatment effect (CATE) compared to later surrogates. This suggests that early surrogates are more sensitive to different treatments, potentially making them superior indicators of the most beneficial treatment compared to the final outcome. In other words, in this scenario, earlier surrogates have the joint advantage of both timeliness and improved sensitivity for detecting treatment effect heterogeneity.
Figure 1: Conditional Average Treatment Effect (CATE) functions for Scenario 1 (left), in which later surrogates have higher utility and Scenario 2 (right) in which earlier surrogates have higher utility. Each curve represents the CATE function of a candidate surrogate, ranging from $Y_{1}$ to $Y_{5}$.
For each design, we run experiments for $T=50$ time points, enrolling 50 subjects at every time point. We set the maximal probability of assigning the superior treatment to be $90\%$. We compare three designs with varying strategies to assign treatment randomization probability: (a) simple sequential RCT, which allocates two treatments with equal probability; (b) adaptive designs with randomization probabilities updated in response to an a priori fixed choice of either a surrogate outcome $Y_{k}$ ($k$ = 1, …, 4) or the final outcome $Y_{K}$ ($K=5$); (c) the proposed adaptive design with Online Superlearner used to adaptively select which surrogate outcome (among candidate surrogates $Y_{1},\cdots,Y_{5}$) is used to update treatment randomization probabilities at each time point.
5.2 Surrogate Evaluation and TMLE Performance in Experiment
We first evaluate if our adaptive design with Online Superlearner can consistently identify the oracle surrogate, which we define as the candidate with the highest true $\Psi_{t,k}$ at time $t$ among all $Y_{k}$ (in other words, the surrogate that if used would result in the optimal expected outcomes for all persons in the trial). The oracle surrogate is the surrogate that results in the adaptive experiment updating treatment randomization probabilities towards true optimal individualized treatments such that benefit for participants is maximized.
Table 1 presents the average of true $\Psi_{t,k}$ values at illustrative time points 11, 21, 31, 41, and 50 for Scenarios 1 and 2 across 500 Monte Carlo simulations. Additionally, $\Psi_{t,RCT}$, the expected outcome of $Y_{5}$ under a non-adaptive simple randomized controlled trial, is reported for both scenarios. The table highlights the expected reward for subjects enrolled up to time $t-K$, under the adaptive design defined by the oracle surrogate, denoted as $\Psi_{t,k_{t}^{*}}$, in bold at each respective time point $t$.
In Scenario 1, the oracle surrogate gradually shifts towards later surrogates as time elapses and more data from these surrogates become available to guide the adaptation of treatment randomization probabilities. Initially, surrogates $Y_{3}$ and $Y_{4}$ have the highest value at times 11 and 21, reflecting the balance they achieve between earlier availability and degree to which they reflect the true optimal individualized treatment assignment. While these surrogates provide benefit in the short term, they do not consistently predict the correct optimal treatment for the primary outcome across all subgroups, leading them to be eventually outperformed by designs that utilize the true primary outcome $Y_{5}$ as additional follow-up time is accrued. This is expected, as the ultimate assessment of a surrogate’s effectiveness as follow-up time extends hinges on its ability to predict the optimal personalized treatment for the primary outcome.
In Scenario 2, as expected, surrogate $Y_{1}$ remains the most useful (i.e. is the oracle surrogate) throughout, due to its early availability, and the fact that it correctly identifies the true optimal individualized treatment for the long-term primary outcome. In constrast, later surrogates $Y_{4}$ and $Y_{5}$ have lower scores. This disparity can be attributed to the fact that while $Y_{4}$ and $Y_{5}$ will eventually serve equally well to identify the optimal personalized treatment as time since trial initiation goes to infinity, their utility is limited under finite trial durations simply because they have been observed for fewer participants.
Table 1: True target estimand $\Psi_{t,k}$ had participants enrolled from $1,\cdots,t-k$ were adaptively treated by a candidate surrogate $Y_{k}$ in Scenarios 1 and 2 ($t=11,21,31,41,50$). Expected outcomes under oracle surrogate shown in bold.
(a) Scenario 1 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 0 -0.010 0.052 0.077 0.065 0.033 21 0 -0.006 0.082 0.144 0.174 0.170 31 0 -0.003 0.089 0.160 0.201 0.206 41 0 -0.001 0.092 0.168 0.214 0.223 50 0 0.001 0.094 0.172 0.221 0.233
(b) Scenario 2 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 0 0.073 0.056 0.039 0.018 0.004 21 0 0.086 0.080 0.072 0.059 0.037 31 0 0.089 0.085 0.081 0.071 0.051 41 0 0.091 0.088 0.084 0.076 0.058 50 0 0.092 0.089 0.086 0.080 0.063
We report the bias, variance, and 95% confidence interval coverage probabilities for the TMLE estimators of $\Psi_{t,RCT}$ and $\Psi_{t,k}$ at time 11, 21, 31, 41, and 50 across two scenarios in Monte Carlo simulations (see Table 4, 5, 6 in Appendix H). The estimators exhibit small bias and low variance, with 95% confidence intervals consistently achieving the nominal coverage rate. These results underpin the reliability of subsequent analyses that calculate the selection frequency of the oracle surrogate within the Online Superlearner.
Figure 2 tracks the frequency with which each candidate adaptive design is selected by our proposed Online Superlearner at each time point $t$. Without prior knowledge of the relationships between various outcomes, the Online Superlearner is able to identify the oracle surrogate outcome for the adaptive design in each setting.
Recall that in Scenario 1, later surrogate outcomes serve as better proxies for the final outcome when determining optimal individualized treatments. The left part of Figure 2 shows that as the experiment progresses, with more observations of the final outcomes available, the Online Superlearner increasingly favors adaptive designs that correspond to later outcomes, ultimately selecting designs related to $Y_{4}$ and $Y_{5}$ with the highest frequency.
Conversely, in Scenario 2, the earliest surrogate outcome $Y_{1}$ emerges as the oracle surrogate, accurately pinpointing the optimal individualized treatment and presenting the highest $\Psi_{t,k}$ among its counterparts. As shown in the right part of Figure 2, the Online Superlearner, without a priori guidance, predominantly opts for adaptive designs that respond to early surrogates, selecting the design corresponding to $Y_{1}$ most frequently, followed by $Y_{2}$ and $Y_{3}$. In contrast, $Y_{4}$ and $Y_{5}$ are rarely chosen.
Figure 2: Frequency of candidate adaptive design selection by the Online Superlearner at each time point, differentiated by outcomes $Y_{k}$ ($k\in\{1,2,3,4,5\}$). The left panel shows Scenario 1 and the right panel Scenario 2. Each color presents a candidate surrogate outcome / final outcome from $Y_{1}$ to $Y_{5}$.The height of each bar indicates the frequency with which $\hat{\psi}_{t,k}$ has the highest lower bound of the confidence interval relative to the other $\hat{\psi}_{t,k^{\prime}}$($k^{\prime}\neq k$).
We also compute the Monte Carlo average of the expected counterfactual final outcome $Y_{5}$ under each candidate adaptive design (see Table 3 in Appendix H). This corresponds to the “ideal target estimand” discussed in Section 2.3 (denoted by $\tilde{\Psi}_{t,k}:=E_{C_{1}^{k},\cdots,C_{n_{t}}^{k}}\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}E_{g_{k,i}\left(\cdot|C_{i}^{k}\right)}\left[Y_{i,K}^{*}\right]$, where $C_{i}^{k}$ is the summary measure for the $i$-th participant in an adaptive design that uses a fixed choice of surrogate $Y_{k}$ to update treatment randomization probability), representing the marginal version of our context-specific target estimand $\Psi_{t,k}$, which is the primary focus of this paper. We observe that our proposed estimand $\Psi_{t,k}$ closely approximates $\tilde{\Psi}_{t,k}$, serving as a fast solution to evaluate the expected outcome under “counterfactual adaptive designs” that update treatment randomization probabilities in response to surrogate outcomes $Y_{k}$, without the need to actually implement that design or estimate that with complex g-computation formula as discussed. The small discrepancy between $\tilde{\Psi}_{t,k}$ and $\Psi_{t,k}$, along with TMLE’s precise estimation and robust inference on $\Psi_{t,k}$, suggests that our method can effectively evaluate the utility of surrogates and optimally choose between candidate surrogates in the adaptive design context.
5.3 Performance of Benefiting Participants in Different Designs
We use $d_{0,K}:\mathcal{W}\to\{0,1\}$ to denote the optimal individualized treatment rule for $Y_{K}$, which is defined as $d_{0,K}(W):=I(\bar{Q}_{0,K}(1,W)-\bar{Q}_{0,K}(0,W)>0)$. This rule assigns treatment based on whether the expected outcome under treatment exceeds that under control, given the baseline covariates of each individual. To quantify the effectiveness of treatment assignments of each design, we define probability (%) of not assigning optimal individualized treatments as $p^{\textit{non-opt}}_{t}:=\frac{1}{E(t)}\sum_{i:t_{i}=t}I\left(A_{i}\neq d_{0,K}(W_{i})\right)$, where $E(t)$ is the number of units enrolled at time $t$. Similarly, $p_{t,RCT}$ and $p_{t,SL}$ denote this probability for the non-adaptive RCT and our proposed adaptive design utilizing the Online Superlearner, respectively. Table 7 (Appendix H) compares these suboptimal treatment assignment probabilities across different designs. We also define the average regret for participants enrolled at time $t$ within an adaptive design as $R_{t}=\frac{1}{E(t)}\sum_{i:t_{i}=t}I\left(A_{i}\neq d_{0,K}(W_{i})\right)\times|\bar{Q}_{0,K}(1,W)-\bar{Q}_{0,K}(0,W)|.$ Here, each participant $i$ contributes the absolute difference in expected outcomes between the optimal and actual treatments. A lower $R_{t}$ indicates a better design that brings benefit in $Y_{K}$ for participants enrolled at time t. We denote $R_{t,k}$ as the average regret within an adaptive design responsive to $Y_{k}$ at time $t$, while $R_{t,\text{RCT}}$ and $R_{t,\text{SL}}$ represent the average regret for the RCT design and our adaptive design with the Online Superlearner, respectively. Table 8 (Appendix H) presents the average regret of different designs across 500 Monte Carlo simulations at times $t=11,21,31,41,$ and $50$. Complementing these, Figure 4 in Appendix H illustrates the regret trajectories with line plots, where the shaded regions represent twice the standard error of $R_{t}$ over the Monte Carlo iterations.
In Scenario 1, where later surrogates are superior for identifying the optimal individualized treatment, the adaptive design responsive to $Y_{5}$, as expected, consistently demonstrates the lowest regret across the experiment. In contrast, in the second scenario, the regret within the adaptive design that responds to $Y_{1}$ remains consistently the lowest. There is a noticeable trend that the adaptive design in response to the final outcome $Y_{5}$ exhibits a slower convergence toward the regret of the oracle surrogate $Y_{1}$, with a substantial discrepancy in regret that persists even at $t=50$.
Of course, in practice, the optimal choice of surrogate would not be known a priori. This motivates the use of an Online Superlearner to data-adaptively select the optimal surrogate and corresponding adaptive design. Across both scenarios, the adaptive design incorporating the Online Superlearner demonstrates stable efficacy in approximating the minimal regret achievable compared to other designs. In our simulations, its regret ranks as the second lowest starting at $t=31$ in Scenario 1, and matches the lowest regret trajectory of $Y_{1}$ following $t=21$ in Scenario 2. The robust performance of our proposed adaptive design can be attributed to the evaluation mechanism embedded within the Online Superlearner, which continuously assesses various surrogate outcomes, strategically selects the most advantageous surrogate that maximizes benefits, and utilizes it in subsequent stages.
5.4 Post-Experiment Inference for Other Causal Estimands
In Table 2, we present the statistical inference for the additional causal estimands introduced in Section 3.5 using the TMLE approach. For the mean of final outcome $Y_{5}$ under the estimated optimal treatment rule for $Y_{5}$, we employ 2-fold CV-TMLE to mitigate finite sample bias. Across all estimands examined, both bias and variance of our estimates remain consistently low, with coverage probabilities of confidence intervals closely aligning with the nominal 95%. These results highlight the robustness and accuracy of TMLE estimators for non-i.i.d. adaptive design data, validating our methods for estimating treatment effects and evaluating dynamic treatment rules even when the data arises from adaptive designs focused on increasing participant benefits and selecting the best surrogates during the experiment, rather than prioritizing inference on these causal estimands.
Table 2: Performance of TMLE estimators for other causal estimands at the end of experiments in Scenarios 1 and 2.
(a) Scenario 1 Estimand Notation Truth Bias (×$\times 10^{-3}$) Var (×$\times 10^{-3}$) Coverage (%) Average Treatment Effect on Final Outcome $Y_{5}$ $\psi_{n,5,T}^{\text{ATE}}$ 0.470 1.60 3.83 95.2 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{1}$ $\psi_{d^{*}_{n,1,T}}$ 0.015 0.92 2.28 95.2 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{2}$ $\psi_{d^{*}_{n,2,T}}$ 0.130 0.55 1.85 94.8 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{3}$ $\psi_{d^{*}_{n,3,T}}$ 0.237 -0.97 1.50 95.4 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{4}$ $\psi_{d^{*}_{n,4,T}}$ 0.313 -0.89 0.97 95.0 Expected Final Outcome $Y_{5}$ under OTR of $Y_{5}$ $\psi_{d^{*}_{n,5,T}}$ 0.338 -2.00 0.87 94.6
(b) Scenario 2 Estimand Notation Truth Bias (×$\times 10^{-3}$) Var (×$\times 10^{-3}$) Coverage (%) Average Treatment Effect on Final Outcome $Y_{5}$ $\psi_{n,5,T}^{\text{ATE}}$ 0.062 0.04 3.89 95.4 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{1}$ $\psi_{d^{*}_{n,1,T}}$ 0.120 1.04 0.49 95.2 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{2}$ $\psi_{d^{*}_{n,2,T}}$ 0.120 1.12 0.50 94.8 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{3}$ $\psi_{d^{*}_{n,3,T}}$ 0.120 0.90 0.51 95.4 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{4}$ $\psi_{d^{*}_{n,4,T}}$ 0.119 0.77 0.51 95.0 Expected Final Outcome $Y_{5}$ under OTR of $Y_{5}$ $\psi_{d^{*}_{n,5,T}}$ 0.106 1.58 1.32 93.0
6 Application
6.1 ADAPT-R Trial
To further illustrate how our adaptive design and TMLE methodologies can be applied in practice, we present an extended simulation study based on real trial data. Our study is inspired by the Adaptive Strategies for Preventing and Treating Lapses of Retention in Human Immunodeficiency Virus (HIV) Care (ADAPT-R) trial (Geng et al.,, 2023), a sequential multiple assignment randomized (SMART) trial designed to evaluate adaptive strategies for improving retention in care among individuals initiating HIV treatment.
The ADAPT-R trial contains a two-phase design. In Phase 1, participants were randomized to the standard of care (SOC), Short Message Service (SMS) reminders, or conditional cash transfers (CCT). Patients who experienced a retention lapse within the first year, defined as missing a clinic visit by $\geq 14$ days, were re-randomized in Phase 2 to a re-engagement strategy: standard-of-care outreach (SOC), combined SMS and CCT (SMS+CCT), or peer navigation (Navigator). We demonstrate our proposed CARA experiments using this real-trial data, showing how it utilizes effective surrogates to benefit participants and evaluating TMLE estimators for causal estimands during and after adaptive experiments.
6.2 Study Setup
Our simulation study focuses on Phase 2 of the ADAPT-R trial, in which patients who experienced a lapse in care within the first year were re-randomized to a re-engagement strategy. We specifically compare the effectiveness of Navigator and SMS+CCT on patient retention over time, as quantified by the proportion of time in care. We discretize the study timeline into 50-day stages and define outcomes as the proportion of time in care across five windows, from the first day of receiving the re-engagement strategy to five specific time points: 50, 100, 150, 200, and 250 days following the initiation of the re-engagement strategy. These outcomes are labeled as $Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$, and $Y_{5}$, respectively, with $Y_{5}$ (proportion of time in care in 250 days) the primary outcome, and $Y_{1}$ through $Y_{4}$ serve as surrogate outcomes. We use data from 215 patients who lapsed after an initial engagement strategy in Year 1 and were assigned either Navigator or SMS+CCT as treatment in Phase 2. Based on these data, we fit a generalized linear model with interaction terms to serve as the data-generating mechanism. The baseline covariates include the Phase 1 treatment assignment (SOC, SMS, or CCT) and the time to lapse after Phase 1 treatment. The CATE functions of different surrogates are presented in Figure 5 in Appendix I. For patients initially assigned to SMS, optimal personalized treatment varies with time-to-lapse after initial treatment, showing different patterns across outcomes. For earlier outcomes ($Y_{1}$ through $Y_{3}$), the CATE suggests that Navigator is preferable with shorter time-to-lapse, but the benefit wanes over time, shifting the optimal treatment to SMS+CCT. In contrast, for later outcomes ($Y_{4}$ and $Y_{5}$), SMS+CCT is preferred for short time-to-lapse, with their effectiveness decreasing over time, making Navigator the optimal choice for longer time-to-lapse. For patients who started with SOC, the treatment effects are minimal: CATE remains close to zero over the entire range of time-to-lapse periods, indicating small differences between Navigator and SMS+CCT. For patients with CCT as their initial strategy, Navigator is preferred for patients with short time-to-lapse periods, showing a positive CATE. However, as time-to-lapse increases, SMS+CCT becomes increasingly favored, with the CATE decreasing and turning negative.
We implement a series of CARA designs, with configurations as outlined in Section 5, including a simple sequential RCT, five adaptive designs that use a fixed choice from $Y_{1}$ through $Y_{5}$, and the proposed adaptive design that employs an Online Superlearner to select the best surrogate over time. Each experiment replicates the patient enrollment schedule in the original trial, ensuring that our simulations capture the timing and dynamics of patient participation. We use the fitted model to generate outcomes based on patients’ baseline covariates, the treatments assigned under the ongoing adaptive design, incorporating noise terms from a normal distribution with a zero mean and a standard deviation based on the model’s standard error estimates. We then simulate each design 500 times to assess the performance of TMLE estimators and different adaptive designs.
6.3 Results
We report the frequency of surrogates chosen by the proposed adaptive design with Online Superlearner in the left plot of Figure 3. Early surrogates (from $Y_{1}$ to $Y_{3}$) are predominantly selected by Superlearner, with $Y_{1}$ and $Y_{2}$ showing increasing selection frequency over time. The right plot of Figure 3 shows the expected final outcome $Y_{5}$ of different adaptive designs over time. Adaptive designs leveraging surrogate information outperform non-adaptive designs consistently. Notably, the adaptive design using the earliest surrogate $Y_{1}$ achieves the highest expected outcome, while later surrogates perform worse. The proposed Online Superlearner design, which dynamically evaluates and selects surrogates, outperforms adaptive designs based on later surrogates ($Y_{4}$ and $Y_{5}$), demonstrating the effectiveness of dynamic surrogate selection over relying on a single outcome when the optimal choice is unknown.
We further assess TMLE’s performance in estimating estimands that evaluate surrogates utilities during the experiment (Table 10) and other causal estimands other estimating causal after experimentation (Table 11). TMLE achieves accurate estimates of the primary causal estimand $\Psi_{t,k}$ ($k=1,\cdots,K$), average treatment effects, and expected final outcome under optimal treatment rules that maximize different surrogates outcomes. All results show consistently nominal coverage probabilities across all phases, highlighting TMLE’s robust estimation capability in adaptive experimental settings.
Figure 3: The left plot illustrates the frequency of candidate adaptive design selections by the Online Superlearner at each time point in the ADAPT-R trial simulation, differentiated by outcomes $Y_{k}$ ($k\in{1,2,3,4,5}$). Each color represents a candidate surrogate or final outcome ($Y_{1}$ to $Y_{5}$). The height of each bar indicates the frequency with which $\hat{\psi}_{t,k}$ achieves the highest lower bound of the confidence interval compared to other $\hat{\psi}_{t,k’}$ ($k’\neq k$). The right plot shows the expected final outcome $Y_{5}$ of different adaptive designs over time.
7 Discussion
In summary, this manuscript contributes to the intersection of surrogate outcomes and adaptive designs in statistical research. We propose a novel target estimand to evaluate surrogate outcomes in the context of sequential adaptive designs. This is defined as the mean of the primary outcome of interest under a counterfactual CARA design in which each participant was randomized based on an estimates optimal individualized treatment rule learned using a given candidate surrogate outcome. This estimand provides a single integrated performance metric that incorporates both how well treatment effects on the candidate surrogate accurately reflect the treatment assignment that will optimize the primary outcome, and also the extent to which the surrogate can benefit more participants at earlier stages through earlier shifts in treatment randomization probability. The proposed estimand thus provides a more comprehensive evaluation of surrogates’ benefits than alternative approaches. We further propose an adaptive design based on an Online Superlearner to allow learning over the course of the study which of a set of candidate surrogates is most useful, and dynamically adapting subsequent randomization probabilities accordingly.
Beyond evaluation of surrogate outcomes, this framework is widely applicable as an approach to evaluate the opportunities and costs of a wide range of candidate adaptive designs, characterized by alternative approaches to adapting randomization probabilities and/or other features of the design. It provides a counterfactual reasoning approach to offer insights into how outcomes might differ under alternative design choices and converge over time, based on incoming data, to the best design for a given sequential decision-making process. We introduce a Targeted Maximum Likelihood Estimation (TMLE) approach to estimate our proposed target estimand from non-i.i.d adpative design data, either during or after the adaptive experiment, obtaining a consistent and asymptotically normal estimator based on Martingale CLT, fundamental maximal inequalities for martingale processes, and the recent theoretical results of Highly Adaptive Lasso. This approach is broadly applicable for constructing robust estimators and providing statistical inference for a general class of causal estimands based on sequentially dependent data collected in multiple setups, including, but not limited to, adaptive sequential designs and time series with temporal dependence.
While we focus on context-specific target estimation in this work, a comprehensive comparison between this and the marginal parameter, as discussed in the Appendix A, remains an interesting question in sequential decision-making setups. Research on this topic could provide valuable insights into the strengths and limitations of these estimands in answering causal queries, as well as the trade-offs with their statistical properties and computational efficiency in different scenarios. We leave the investigation of this topic for future research.
SUPPLEMENTARY MATERIAL of
“Evaluating and Utilizing Surrogate Outcomes in Covariate-Adjusted Response-Adaptive Designs”
Appendix A Additional Remarks on Target Estimands
Remark 1 on two types of general target estimands in Section 2.3. While the cumulative impact over $\mathcal{\bar{C}}^{*}(n)$ and $\mathbf{g}_{*,1:n}(\mathcal{\bar{C}}^{*}(n))$ may appear unpredictable, such concerns are mitigated under a well-designed CARA framework, which makes our target estimand $\Psi_{\mathbf{g}_{*,1:n}\left(\mathcal{\bar{C}}(n)\right)}$ a viable alternative estimand to circumvent the identification and estimation complexities by the $g$-formula. Specifically, in Section 3, we discuss the CARA framework where each $C_{i}$ only depends on CATE functions estimated from previous data and the surrogate selected by Online Superlearner. In this setting, any change in early context primarily affects the proportion of treatment and control observed in subsequent stages, without altering the true CATE functions to be estimated. By applying a doubly-robust estimator for CATE functions using pseudo-outcomes (van der Laan,, 2006; Luedtke and van der Laan, 2016b, ), we expect such impact will be mild on future contexts and the choice of surrogate in the Online Superlearner in our setup, thanks to the stationarity in the outcome models and our knowledge of true treatment probabilities in the adaptive design. We refer readers to Section 9.1 in Malenica et al., (2021) for a detailed discussion about the statistical and computational advantages of context-specific estimands compared to marginal target estimands.
Remark 2 on the proposed target estimand for evaluating surrogates in Section 3.2. Our target estimand evaluates the utility of surrogates by their early availability, as well as their predictability and sensitivity in indicating effective treatments, defined in the following sense. We define a trajectory of stochastic treatment rules $\tilde{g}^{*}_{k,1:t}:=(\tilde{g}^{*}_{k,1},\cdots,\tilde{g}^{*}_{k,t})$, with each rule $\tilde{g}^{*}_{k,t^{\prime}}$ generating personalized stochastic rules $g^{*}_{k,i}$ for each participant $i$ enrolled at time $t^{\prime}$. The rule $g^{*}_{k,i}$ is specifically tailored to maximize the surrogate outcome $Y_{k}$ based on their baseline covariates. We note that at each time point $t^{\prime}$, $\tilde{g}^{*}_{k,t^{\prime}}$ only depends on historical data ¯$\overline{\mathcal{H}}(t^{\prime})$. Therefore, $\tilde{g}^{*}_{k,t^{\prime}}$ can not deviate from a static design with equal randomization probabilities during the initial $k$ time points, as observation of $Y_{k}$ is not yet available. Once $Y_{k}$ is available at $t^{\prime}=k+1$, $\tilde{g}^{*}_{k,t^{\prime}}$ starts moving towards its optimal treatment rule $d_{0,k}(W):=I\left(\bar{Q}_{0,k}(1,W)-\bar{Q}_{0,k}(0,W)>0\right)$ over time. If $d_{0,k}$ closely aligns with the true optimal rule for the primary outcome $d_{0,K}$, then the surrogate $Y_{k}$ demonstrates the advantage of its early availability in informing which treatment is better in the context of sequential decision making, compared to later outcomes. As is shown in Section 3.1, the progression of $\tilde{g}^{*}_{k,t^{\prime}}$ from a static design with equal randomization probabilities towards the optimal treatment rule $d_{0,k}$ depends on the magnitude of treatment superiority (as measured by the point estimate of CATE) and its uncertainty (as reflected by estimated standard error of CATE). Therefore, if a surrogate $Y_{k}$ exhibits good alignment with the primary outcome in terms of optimal treatment rule and presents a strong signal of individualized optimal treatment over other surrogates, this surrogate will be preferred and given a higher $\Psi_{t,k}$. Of course, even if a surrogate demonstrates early availability and sensitivity, it will be assigned a lower $\Psi_{t,k}$ value if it fails to predict optimal personalized treatments. This is because the ultimate evaluation of a surrogate’s performance is based on the primary outcome, underscoring the critical role of predictability in evaluating surrogate outcomes in adaptive designs. In summary, our proposed target estimand quantifies the utility of a candidate surrogate and thus provides a summary measure that incorporates these trade-offs.
Appendix B The Highly Adaptive Lasso Estimator
The Highly Adaptive Lasso (HAL) (van der Laan,, 2015; Benkeser and van der Laan,, 2016; van der Laan,, 2023) is a nonparametric estimator for regression functions that does not require local smoothness of the true regression function. In short, the HAL estimator is constructed from a set of data-adaptive Càdlàg basis functions (i.e., right-hand continuous functions with left-hand limits) and is represented as the minimizer of an empirical risk criterion subject to a $L_{1}$ constraint on the sectional variation of the target regression function. In this paper, we use HAL to estimate Conditional Average Treatment Effect (CATE) functions (see Appendix C) and provide a corollary for proving equicontinuity condition of martingale process based on recent theoretical results of higher order spline HAL-MLEs in (van der Laan,, 2023) (see Appendix E, F).
We first introduce $D^{(0)}([0,1]^{d})$, a class of $d$-variate cadlag functions defined on a unit cube $[0,1]^{d}$ with bounded sectional variation norm (Gill et al.,, 1995) and the Highly Adaptive Lasso Estimator for $D^{(0)}([0,1]^{d})$ (van der Laan,, 2015; Benkeser and van der Laan,, 2016; Bibaut and van der Laan,, 2019). Then we will introduce the recent result of van der Laan, (2023) that introduces $D^{(m)}([0,1]^{d})$ $(m=1,2,\cdots)$, a class of $m$-th order smoothness functions with bounded sectional variation norm on up to their $k$-th-order derivatives and the higher order spline HAL-MLEs to estimate them.
According to Gill et al., (1995); van der Laan, (2015); Benkeser and van der Laan, (2016), any function $Q\in D^{(0)}([0,1]^{d})$ can be represented as
[…]
with basis function $\phi_{u}^{0}(x):=I(u\leq x)=\prod_{j=1}^{d}I\left(u_{j}\leq x_{j}\right)$ a tensor product of zero-order splines. For every non-empty $s\subset\{1,\ldots,d\}$, we can define $E_{s}=\left\{(x(s),0(-s)):x\in[0,1]^{d}\right\}$ as a $s$-specific zero edge of $[0,1]^{d}$, which sets the coordinates in the complement of subset $s$ equal to 0. As a convention, we include the empty subset $s=\emptyset$ in which case $E_{\emptyset}=\{0\}$. Since $[0,1]^{d}$ can be partitioned into the singleton $\{0\}$ ($s=\emptyset$), $s$-specific left-edges $E_{s}$ ($s\subsetneqq\{1,\ldots,d\}$, $s\neq\emptyset$), and the full-dimensional inner set $(0,1]^{d}$ ($s=\{1,\ldots,d\}$), one can represent $dQ(u)$ as the sum measure over these disjoint complementary supports $dQ(u)=\sum_{s\subset\{1,\ldots,d\}}I(u\in E_{s})dQ(u(s),0(-s))$. For $s=\emptyset$, $E_{s}=\{0\}$, $dQ(u(s),0(-s))=dQ(\{0\})=Q(0)$; for a non-empty $s\subsetneqq\{1,\ldots,d\}$, $dQ(u(s),0(-s))$ is the measure $Q(du(s),0(-s))$ implied by its $s$-specific section $Q_{s}:x(s)\to Q(x(s),0(-s))$ on the zero edge $E_{s}$; for $s=\{1,\ldots,d\}$, $dQ(u(s),0(-s))$ represents the measure for $u$ in the full-dimensional inner cube $(0,1]^{d}$. Therefore, any cadlag function $Q\in D^{(0)}([0,1]^{d})$ can also be written as
[…]
The sectional variation norm of $Q$ is defined by the sum over $s$ of the variation norms of the measures on every $s$-specific section:
$\|Q\|_{v}^{*}:=\int_{[0,1]^{d}}|dQ(u)|=|Q(0)|+\sum_{s\subset\{1,\ldots,d\},s\neq\emptyset}\int_{(0(s),1(s)]}|Q(du(s),0(-s))|.$
We use $[d]$ to denote the full set $\{1,\ldots,d\}$ and correspondingly define $\mathcal{S}^{0}_{[d]}$ as the set of all subsets $s\subset[d]$, including the empty subset. For each subset $s\subset\{1,\ldots,d\}$, we define $\mathcal{R}^{0}(s)=(0(s),1(s)]$ on the $s$-specific zero edge $E_{s}$. Let [$\mathcal{R}^{0}_{[d]}=\left\{(s,u_{s}):s\neq\emptyset,s\in\mathcal{S}^{0}_{[d]},u_{s}\in\mathcal{R}^{0}(s)\right\}\cup\left\{\left(\{0\},0\right):s=\emptyset\right\}$ be the set of all subsets $s\subset[d]$ and all knot points $u_{s}$ on $s$-specific zero edges, which represents all basis functions needed to span $D^{(0)}([0,1]^{d})$. Therefore, one can represent $D^{(0)}([0,1]^{d})$ by
[…]
One can define [$\mathcal{R}^{0,*}_{[d]}=\left\{(s,u_{s}):s\in\mathcal{S}^{0,*}_{[d]},u_{s}\in\mathcal{R}^{0,*}(s)\right\}$, where $\mathcal{S}^{0,*}_{[d]}\subset\mathcal{S}^{0}_{[d]}$ and $\mathcal{R}^{0,*}(s)\subset\mathcal{R}^{0}(s)$ for every $s\in[d]$. Since [$\mathcal{R}^{0,*}_{[d]}$ is a subset of [$\mathcal{R}^{0}_{[d]}$, one can define the following submodel of $D^{(0)}([0,1]^{d})$ based on [$\mathcal{R}^{0,*}_{[d]}$:
[…]
One can also specify a submodel whose sectional variation norm $\|Q\|_{v}^{*}$ is upper bounded by $C^{u}$:
[…]
Suppose we observe $n$ i.i.d samples $(x_{i},y_{i})$ ($i=1,\cdots,n$) with $x_{i}=(x_{i,1},\cdots,x_{i,d})\in\mathbb{R}^{d}$, $y_{i}\in\mathbb{R}$. van der Laan, (2015); Benkeser and van der Laan, (2016) proposed to use discrete measures that put mass on the observed knot points $\mathcal{R}_{n}^{0,*}(s)=(x_{i}(s),i\in[n])\subset\mathcal{R}^{0,*}(s)$ on all $s$-specific edges ($s\subset[d]$) to construct a submodel $D^{(0)}(\mathcal{R}^{0,n}_{[d]})$ to approximate the representation of $Q(x)$:
$D^{(0)}(\mathcal{R}^{0,n}_{[d]}):=\left\{\sum_{s\in\mathcal{S}^{0,*}_{[d]}}\sum_{u_{s}\in\mathcal{R}_{n}^{0,*}(s)}\phi_{u_{s}}^{0}(\cdot)dQ(u_{s}):Q\in D^{(0)}([0,1]^{d})\right\},$
where [$\mathcal{R}^{0,n}_{[d]}:=\left\{(s,u_{s}):s\neq\emptyset,s\in\mathcal{S}^{0,*}_{[d]},u_{s}\in\mathcal{R}_{n}^{0,*}(s)\right\}\cup\left\{\left(\{0\},0\right):s=\emptyset\right\}$. Similarly, one can consider a submodel with bounded total variation norm $C^{u}$:
$D^{(0)}_{C^{u}}(\mathcal{R}^{0,n}_{[d]}):=\left\{\sum_{s\in\mathcal{S}^{0,*}_{[d]}}\sum_{u_{s}\in\mathcal{R}_{n}^{0,*}(s)}\phi_{u_{s}}^{0}(\cdot)dQ(u_{s}):Q\in D^{(0)}([0,1]^{d}),\|Q\|_{v}^{*}<C^{u}\right\}.$
Accordingly, a zero-order HAL-MLE can be defined by solving the following loss-based minimization problem:
$Q_{n}=\mathop{\mathrm{argmin}}_{Q\in D^{(0)}_{C_{n}^{u}}(\mathcal{R}^{0,n}_{[d]})}\mathbb{P}_{n}L(Q),$
with $L(\cdot)$ an appropriately defined loss function and $C_{n}^{u}$ a data-adaptive cross-validation or undersmoothed selector of bounded sectional variation norm for the working model. Define |[$\big{|}\mathcal{R}^{0,n}_{[d]}\big{|}$ the cardinality of [$\mathcal{R}^{0,n}_{[d]}$ and $\mathcal{B}_{n}:=\left\{\beta:\beta\in\mathbb{R}^{\big{|}\mathcal{R}^{0,n}_{[d]}\big{|}};\beta=\left(\beta_{s,u_{s}}\right),(s,u_{s})\in\mathcal{R}_{n}^{0,*}(s)\right\}$. The construction of HAL-MLE is equivalent to solving
$\beta_{n}=\mathop{\mathrm{argmin}}_{\beta:\beta\in\mathcal{B}_{n},\|\beta\|_{1}\leq C_{n}^{u}}\mathbb{P}_{n}L\left(\sum_{s\in\mathcal{S}^{0,*}_{[d]}}\sum_{u\in\mathcal{R}_{n}^{0,*}(s)}\phi_{u_{s}}^{0}(\cdot)\beta_{s,u_{s}}\right),$
thus HAL-MLE can be expressed by $Q_{n}(\cdot)=\sum_{s\in\mathcal{S}^{0,*}_{[d]}}\sum_{u_{s}\in\mathcal{R}_{n}^{0,*}(s)}\phi_{u_{s}}^{0}(\cdot)\beta_{s,u_{s}}$. Bibaut and van der Laan, (2019) shows the $L_{2}$ rate of convergence of $O(n^{-\frac{1}{3}})$ up till a $\log$-n factor for zero-order HAL-MLE.
[…]
Appendix C Estimation of Conditional Average Treatment Effect
A key procedure in our adaptive design is estimating the Conditional Average Treatment Effect (CATE). By estimating the CATE function, the experimenter can identify the optimal treatment rule that maximizes the expected reward based on the expected outcome conditional on baseline covariates, and update treatment randomization probabilities in adaptive design as in Subsection 3.1.
We propose using the Highly Adaptive Lasso (HAL) estimator to estimate the CATE function. As introduced in Appendix B, the HAL estimator is a nonparametric regression method that can estimate a rich class of cadlag functions with a remarkable rate of convergence without relying on local smoothness assumptions. Furthermore, we can directly apply the delta-method to obtain a Wald-type confidence interval for the CATE. This confidence interval guides the exploration-exploitation trade-off in adaptive designs by adjusting treatment randomization probabilities based on the uncertainty in CATE estimates.
Recall that $\bar{Q}_{0,k}(A,W)=E[Y_{k}|A,W]$ is the true conditional expectation of the $k$-th surrogate outcome, given the treatment $A$ and baseline covariates $W$. For every new participant $i$ enrolled at time $t$, we use $B_{n,k,t}(w_{i})$ to denote an estimate of $B_{0,k}(w_{i})$, the true CATE of $Y_{k}$ given $w_{i}$, the baseline covariates of participant $i$. We use $\tau_{n,k,t}(w_{i})$ to denote the estimated standard error of CATE. Both estimates are based on $\mathcal{H}(i)$, the historical data observed before participant $i$ is enrolled.
For each surrogate outcome $Y_{k}$, we estimate CATE functions by two steps. First, we transform $Y_{k}$ into the pseudo-outcome (van der Laan,, 2006; Luedtke and van der Laan, 2016b, ):
$\eta_{k}(\bar{Q}_{k},g)(O)=\frac{2A-1}{g(A\mid W)}[Y_{k}-\bar{Q}_{k}(A,W)]+\bar{Q}_{k}(1,W)-\bar{Q}_{k}(0,W).$
We note that the pseudo-outcome has the doubly-robust property that $E[\eta_{k}(\bar{Q}_{k},g)(O)\mid W]=B_{0,k}(W)$ if $\bar{Q}_{k}=\bar{Q}_{0,k}$ or $g=g_{0}$. Since $g_{0}$ is known in adaptive design, we only need to estimate $\bar{Q}_{0,k}$ with $\bar{Q}_{n,k}$, then impute $\bar{Q}_{n,k,t}$ and $g_{0}$ to generate the pseudo-outcome $\eta_{k}(\bar{Q}_{n,k,t},g_{0})$.
Second, we estimate the CATE function $B_{0,k}$ by regressing the pseudo-outcome $D_{k}\left(\bar{Q}_{n,k,t},g_{0}\right)$ on baseline covariates $W$ by first-order spline HAL-MLE, which minimizes the empirical risk over the $D^{(1)}([0,1]^{d})$, the first-order smoothness class with its first order derivative being cadlag function with bounded first order sectional variation norm. That is, at time $t$, we estimate the CATE function by solving the following minimization problem:
$B_{n,k,t}=\mathop{\mathrm{argmin}}_{B_{k}\in D^{(1)}_{C_{n}^{u}}(\mathcal{R}^{1,n}_{[d]})}\frac{1}{n_{t,k}}\sum_{j=1}^{n_{t,k}}L_{k}(B_{k})(O_{j}),$
where $L_{k}$ is the loss function defined by $L_{k}(B_{k})(O):=(B_{k}(W)-\eta_{k}(O))^{2}$; $n_{t,k}:=N(t-k)$ is the number of participants with observed $Y_{k}$ at time $t$; [$\mathcal{R}^{1,n}_{[d]}$ is the set of observed knot points from samples (see Appendix B); $C_{n}^{u}$ is an upper bound for the first order sectional variation norm that is data-adaptively selected via cross-validation.
As introduced in Appendix B, $B_{n,k,t}(\cdot)$ can be written as a linear combination of first-order spline basis functions with non-zero coeffcients. To quantify the uncertainty of CATE estimates, we apply the delta method in this data-adaptive HAL working model to make inferences on CATE. We use $\tau_{n,k,t}(w_{i})$, the estimate of the standard error of $B_{n,k,t}(w_{i})$, to construct a Wald-type confidence interval for the CATE function. Both $B_{n,k,t}(w_{i})$ and $\tau_{n,k,t}(w_{i})$ are incorporated into the $h_{m}$ function in Section 3.1 to determine the treatment randomization probability $g_{k,i}^{*}$ for every participant $i$.
Appendix D Details of Our CARA Procedures to Generate Treatment Randomization Probabilities
We define a treatment randomization function $h_{m}:[-1,1]\times\mathbb{R}^{+}\to[0,1]$ as follows:
$h_{\nu}(x,b)=\nu\mathds{1}(x\leq-b)+(1-\nu)\mathds{1}(x\geq b)+\left(\frac{1/2-\nu}{2b^{3}}x^{3}+\frac{1/2-\nu}{2b/3}x+\frac{1}{2}\right)\mathds{1}\left(-b\leq x\leq b\right),$
where $\nu\in[0,0.5]$ is a user-specified constant such that $\nu<|h_{\nu}|<1-\nu$. Let $B_{0,k}:\mathcal{W}\to[-2,2]$ denote the true CATE function $E[Y_{k}\mid 1,W]-E[Y_{k}\mid 0,W]$ of an outcome $Y_{k}$ ($k=1,\cdots,K$). Let $B_{n,k}(w_{i})$ denote an estimate of CATE of $Y_{k}$ based on $w_{i}$, the baseline covariates of unit $i$. We use $\tau_{n,k}(w_{i})$ to denote the estimated standard error of CATE of unit $i$. Both estimates are based on $\mathcal{H}(i)$, the historical data observed before unit $i$ is enrolled. We employ a doubly-robust estimator to estimate CATE functions using pseudo-outcomes (van der Laan,, 2006; Luedtke and van der Laan, 2016b, ) and applying Highly Adaptive Lasso (van der Laan,, 2015, 2017; Benkeser and van der Laan,, 2016; van der Laan,, 2023) to obtain $B_{n,k}$ and $\tau_{n,k}$ (see Appendix C). Users are flexible to define their own methods to estimate them.
Then we use $h_{\nu}$ to map $C_{i}$ to a treatment randomization probability between 0 and 1. We define a three-dimensional, $k$-specific summary measure $C_{i,k}=\left(w_{i},B_{n,k}(w_{i}),\tau_{n,k}(w_{i})\right)$ for every $k\in[K]$, the index for surrogate $Y_{k}$. Taking the union of $C_{i,k}$ across $k\in[K]$, we define $C_{i}:=(C_{i,k})_{k=1}^{K}$ as the complete summary measure before treatment decision for unit $i$. Recall that $g^{*}_{k,i}(a\mid C_{i})$ describes the treatment randomization probability tailored for optimizing a surrogate outcome $Y_{k}$, based on summary measure $C_{i}$. $g^{*}_{k,i}$ is generated by the following equation:
$g^{*}_{k,i}(a|C_{i})=\begin{cases}h_{\nu}\left(B_{n,k}(w_{i}),z_{1-\alpha/2}\cdot\tau_{n,k}(w_{i})\right)&a=1,\\ 1-h_{\nu}\left(B_{n,k}(w_{i}),z_{1-\alpha/2}\cdot\tau_{n,k}(w_{i})\right)&a=0,\end{cases}$
where $z_{1-\alpha/2}$ is the $21-\alpha/$ 2 quantile of standard normal distribution ($\alpha=0.05$).
Our strategy balances exploration and exploitation based on the CATE functions learned from the data. Intuitively, $g^{*}_{k,i}(a|C_{i})$ depends on the confidence interval of the estimated CATE of $Y_{k}$ for participant $i$: $B_{n,k,i}(w_{i})\pm z_{1-\alpha/2}\cdot\tau_{n,k,i}(w_{i})$. The more deviated the confidence interval is from zero, the more likely it is to assign unit $i$ to an optimal individual treatment. Note that $\nu$ is the lower bound of the randomization probability of assigning a sub-optimal treatment, and $1-\nu$ is the upper bound of the probability to assign optimal treatment. If the confidence interval excludes zero, we assign the better treatment with probability $1-\nu$. If no $Y_{k}$ has been observed before the enrollment of unit $i$, we set $g^{*}_{k,i}=0.5$.
Remark 3. We note that in adaptive designs in which randomization probabilities are adapted in response to an outcome $Y_{k}$, all participants would be randomly assigned to different treatment groups with equal probability prior to the first observation of surrogate $Y_{k}$ at $t=k+1$ (Section 3.1). Similarly, in our adaptive design with the Online Superlearner, every treatment decision before $t=K+1$ is randomly made by flipping a fair coin. This is because the surrogate evaluation procedure is only activated after the primary outcome $Y_{K}$ has been observed for the initially enrolled subset of participants. While we will propose in Appendix J a modified adaptive design that is able to initiate surrogate evaluation and the adapt treatment randomization probabilities at an earlier stage, the adaptive designs investigated within this section represent the most conservative approaches ensuring that every adaptation, without exception, is activated only subsequent to the observation of related outcomes within the ongoing experimental environment. Such conservativeness helps maintain reliability in adapting treatment randomization probabilities and evaluating surrogates.
While the optimal choice of a surrogate is not known a priori in practice, our Online Superlearner will identify the most useful surrogate for maximizing reward as more data accumulates. It will then adjust the actual treatment randomization probabilities towards the optimal treatment rule associated with that surrogate. Consequently, our design will converge to a design that improves the efficiency in estimating the mean outcome under the optimal treatment rule of that most useful surrogate.
Appendix E General Theorem for Equicontinuity Condition of a General Martingale Process
In this section, we provide a theorem about the equicontinuity condition of martingale processes in a setup that is more general than our adaptive design.
E.1 Statistical Setup
Suppose we sequentially generate a sequence of data $O_{i}$ from $i=1$ to $N$. Every $O_{i}\in\mathcal{O}_{i}$ is generated from historical data $\bar{O}(i-1):=(O_{1},\cdots,O_{i-1})$. We define a probability triple $(\Omega,\mathcal{F},\mathbf{P})$, where $\mathcal{F}_{i}$, $(i=1,\cdots,N,\cdots)$ are defined as sub-sigma algebras $\mathcal{F}_{1}\subset\cdots\mathcal{F}_{N}\subset\cdots\subset\mathcal{F}$ with $\bar{O}(i)\in\mathcal{O}_{1}\times\cdots\times\mathcal{O}_{i}\subset\mathcal{F}_{i}$. We consider a scenario in which every $O_{i}$ is generated based on $\bar{O}(i-1)$ through a fixed dimensional summary measure $C_{i}\in\mathcal{C}$, where $C_{i}$ is generated by a known function mapping from $\mathcal{O}_{1}\times\cdots\times\mathcal{O}_{i-1}$ to $\mathcal{C}$ for every $i=1,\cdots,N$.
Let $f_{\theta}:\mathcal{O}\times\mathcal{C}\to\mathbb{R}$ be a $\mathcal{F}_{i}$-measurable function, indexed by a class of functions $\Theta$. Define
$M_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}f_{\theta}(O_{i},C_{i})-E\left[f_{\theta}(O_{i},C_{i})\mid\bar{O}(i-1)\right].$
Suppose there is an unknown element $\theta^{*}\in\Theta$ to be estimated by $\theta_{N}\in\Theta$ based on data $\bar{O}(N)$. We provide a general theorem for the equicontinuity condition with respect to the martingale process as follows:
$M_{2,N}(\theta_{N},\theta^{*}):=M_{N}(\theta_{N})-M_{N}(\theta^{*})=o_{P}(N^{-\frac{1}{2}}).$
E.2 Proof of Equicontinuity Condition
We first restate a bracketing entropy based on Definition A.1 in van Handel, (2011) (which was introduced by Definition 8.1 in van de Geer, (1999)) by our notation.
{restatable}
definitionrepDefSeqBracketing(Bracketing entropy for martingales) Consider a stochastic process of the form $\left\{\left(\xi_{\theta,i}\right)_{i=1}^{N}:\theta\in\Theta\right\}$ where $\Theta$ is an index set such that, for every $\theta\in\Theta,i\in[N]$, $\xi_{\theta,i}$ is an $\bar{O}(i)$-measurable random variable for all $i$ and $\theta$. Denote $F:=\bar{O}(N)\in\mathcal{F}$. We say that a collection of random variables of the form :=$\mathcal{B}:=\left\{\left(\Lambda_{i}^{j},\Upsilon_{i}^{j}\right)_{i=1}^{N}:j\in[J]\right\}$ is a $(N,\Theta,F,Z,\epsilon)$-bracketing if
  1. 1.
    for every $i\in[N]$, and $j\in[J],\left(\Lambda_{i}^{j},\Upsilon_{i}^{j}\right)$ is $\bar{O}(i)$-measurable,
  2. 2.
    for every $\theta\in\Theta$, there exists $j\in[J]$, such that, for every $i\in[N],\Lambda_{i}^{j}\leq\xi_{\theta,i}\leq\Upsilon_{i}^{j}$,
  3. 3.
    for every $j\in[J]$,
    $\frac{2Z^{2}}{N}\sum_{i=1}^{N}E\left[\phi\left(\frac{\Upsilon_{i}^{j}-\Lambda_{i}^{j}}{Z}\right)\mid\bar{O}(i-1)\right]\leq\epsilon^{2},$
[…]
Theorem 5 (Theorem A.4 in van Handel, (2011)).
Given $Z>0$, for all $i\geq 0$, define
$M_{\theta,i}=\frac{1}{N}\sum_{l=1}^{i}\left(\xi_{\theta,l}-E\left[\xi_{\theta,l}\mid\mathcal{F}_{l-1}\right]\right),$
$R_{\theta,i}=\frac{1}{N}2Z^{2}\sum_{l=1}^{i}E\left[\phi\left(\frac{|\xi_{\theta,l}|}{Z}\right)\bigg{|}\mathcal{F}_{l-1}\right].$
By Theorem A.4 in van Handel, (2011), for any $N\in\mathbb{N}$, $x>0$, and $0<R<\infty$,
$P\left(\sup_{\theta\in\Theta}I\left(R_{\theta,N}\leq R\right)\max_{i\in[N]}\sqrt{N}M_{\theta,i}\geq 16\Gamma+32\sqrt{Rx}+\frac{16}{\sqrt{N}}Zx\right)\leq 2e^{-x},$
where $\Gamma:=\frac{Z}{\sqrt{N}}\log\mathcal{N}_{[]}\left(N,\Theta,F,{Z},{\sqrt{R}}\right)+4\mathcal{J}_{[]}\left(N,\Theta,F,{Z},{\sqrt{R}}\right)$.
We focus on the process of $\xi_{\theta,i}=f_{\theta}(O_{i},C_{i})-f_{\theta^{*}}(O_{i},C_{i})$ ($i=1,\ldots,n$), with $\theta^{*}\in\Theta,\theta\in\Theta$. As the bracketing entropy condition may be hard to check because of its intricate definition, we relate it to the supremum norm covering entropy, following the approach outlined by van de Geer, (1999) in Definition 8.1 and Chambaz and van der Laan, (2011) in Lemma 7. Our method slightly differs from their approaches in that we relate the bracketing entropy to the supremum norm covering entropy for $\Theta$ instead of $\{f_{\theta}:\theta\in\Theta\}$.
[…]
Assumption 7 (Uniform Boundedness).
$\sup_{\theta\in\Theta}\left\|\theta\right\|_{\infty}<\infty$, $\sup_{\theta\in\Theta}\left\|f_{\theta}\right\|_{\infty}<\infty$.
Assumption 8.
$\sup_{\theta_{1}\in\Theta,\theta_{2}\in\Theta,\theta_{1}\neq\theta_{2}}\frac{\left\|f_{\theta_{1}}-f_{\theta_{2}}\right\|_{\infty}}{\left\|\theta_{1}-\theta_{2}\right\|_{\infty}}<\infty$.
Lemma 1.
Suppose that Assumptions 7 and 8 hold, such that $\left\|\theta\right\|_{\infty}<M$ for some finite constant $M>0$ and $\frac{\left\|f_{\theta_{1}}-f_{\theta_{2}}\right\|_{\infty}}{\left\|\theta_{1}-\theta_{2}\right\|_{\infty}}\leq L$ for some finite constant $L>0$. Then, if $N_{\infty}(\epsilon,\Theta)$ is finite, the following holds:
$\mathcal{N}_{[]}\left(N,\Theta,F,{8LM},{2\sqrt{2}L\epsilon}\right)\leq\mathcal{N}_{\infty}(\epsilon,\Theta),$
and
$\mathcal{J}_{[]}\left(N,\Theta,F,{8LM},{2\sqrt{2}L\epsilon}\right)\leq\mathcal{J}_{\infty}\left({\epsilon},{\Theta}\right).$
Proof.
Suppose $\{\theta_{j}:j\in[J]\}\subset\Theta$ is an $\epsilon$-covering of $\Theta$ in supremum norm such that $\forall\theta\in\Theta$, there exists $j\in[J]$ such that $\theta_{j}\in\Theta$ and $\|\theta_{j}-\theta\|_{\infty}\leq\epsilon$. Denoting $\lambda_{j}=\theta_{j}-\epsilon$ and $v_{j}=\theta_{j}+\epsilon$. With Assumption 7, we can let $\epsilon\leq 2M$ as $\lambda_{j}=\theta_{j}-M$ and $v_{j}=\theta_{j}+M$ always construct a valid pair of brackets. Under Assumption 8, $\left\|f_{\theta_{j}}-f_{\theta}\right\|_{\infty}\leq L\left\|\theta_{j}-\theta\right\|_{\infty}\leq L\epsilon$. Define $\Lambda_{i}^{j}=f_{\theta_{j}}-L\epsilon$ and $\Upsilon_{i}^{j}=f_{\theta_{j}}+L\epsilon$. Then $|\Upsilon_{i}^{j}-\Lambda_{i}^{j}|\leq 2L\epsilon$. Therefore, we have that
$\frac{1}{N}2Z^{2}\sum_{i=1}^{N}E\left[\phi\left(\frac{|\Upsilon_{i}^{j}-\Lambda_{i}^{j}|}{Z}\right)\mid\bar{O}(i-1)\right]\leq 2Z^{2}\phi\left(\frac{2L\epsilon}{Z}\right)\leq 2Z^{2}\sum_{p\geq 2}\frac{1}{p!}\frac{p!}{2}\left(\frac{2L\epsilon}{Z}\right)^{p}=\frac{2L\epsilon}{\frac{1}{2L\epsilon}-\frac{1}{Z}}.$
Let $Z=8LM$. We have that
$\frac{1}{N}2Z^{2}\sum_{i=1}^{N}E\left[\phi\left(\frac{|\Upsilon_{i}^{j}-\Lambda_{i}^{j}|}{Z}\right)\mid\bar{O}(i-1)\right]\leq(2\sqrt{2}L\epsilon)^{2}.$
Therefore,
$\mathcal{N}_{[]}\left(N,\Theta,F,{8LM},{2\sqrt{2}L\epsilon}\right)\leq\mathcal{N}_{\infty}(\epsilon,\Theta),$
and
$\mathcal{J}_{[]}\left(N,\Theta,F,{8LM},{2\sqrt{2}L\epsilon}\right)\leq\mathcal{J}_{\infty}\left({\epsilon},{\Theta}\right).$
Assumption 9 (Convergence of $\theta_{N}$).
Define
$s_{N}\left(\theta,\theta^{*}\right):=\sqrt{\frac{1}{N}\sum_{i=1}^{N}E\left[\phi\left(\frac{f_{\theta}\left(O_{i},C_{i}\right)-f_{\theta^{*}}\left(O_{i},C_{i}\right)}{Z}\right)\mid\bar{O}(i-1)\right]},$
where $Z=8LM$. $\theta_{N}$ is an estimate of $\theta^{*}$ based on $\bar{O}(N)$ such that $s_{N}\left(\theta_{N},\theta^{*}\right)=o_{P}(1)$.
Theorem 6 (Maximal Inequality of Martingale Process).
$\forall x>0,N\in\mathbb{N}$, and $0<r<\infty$. Denote $a_{N}(r,x):=\frac{16Z}{\sqrt{N}}\log\mathcal{N}_{\infty}\left({\frac{r}{2\sqrt{2}L}},{\Theta}\right)+64\mathcal{J}_{\infty}\left({\frac{r}{2\sqrt{2}L}},{\Theta}\right)+32r\sqrt{x}+\frac{16}{\sqrt{N}}Zx$. Suppose that Assumptions 7, 8 and 9 hold. Then,
$P\left(\sup_{\theta^{*}\in\Theta}I\left(s_{N}(\theta_{N},\theta^{*})\leq r\right)\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})\geq a_{N}(r,x)\right)\leq 2e^{-x}.$
Proof.
Recall that
$M_{2,N}(\theta_{N},\theta^{*})=\frac{1}{N}\sum_{i=1}^{N}f_{\theta}\left(O_{i},C_{i}\right)-f_{\theta^{*}}\left(O_{i},C_{i}\right)-E\left[f_{\theta}(O_{i},C_{i})-f_{\theta^{*}}(O_{i},C_{i})\mid\bar{O}(i-1)\right].$
Let $Z=8LM$ and $b_{N}(r,x)=\frac{16Z}{\sqrt{N}}\log\mathcal{N}_{[]}\left(N,\Theta,F,{Z},{r}\right)+64\mathcal{J}_{[]}\left(N,\Theta,F,{Z},{r}\right)+32r\sqrt{x}+\frac{16}{\sqrt{N}}Zx$. We have that
$P\left(\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})I\left(s_{N}\left(\theta_{N},\theta^{*}\right)\leq r\right)\geq a_{N}(r,x)\right)$
$\leq P\left(\sup_{\theta^{*}\in\Theta}I\left(\sigma_{\theta,N}\leq r\right)\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})\geq a_{N}(r,x)\right)$
$\leq P\left(\sup_{\theta^{*}\in\Theta}I\left(\sigma_{\theta,N}\leq r\right)\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})\geq b_{N}(r,x)\right)$
$\leq 2e^{-x},$
where the second inequality is given by Lemma 1 and the third inequality is based on the maximal inequality for the martingale process in Theorem 5.
{restatable}
corollaryrepCoroHighOrderSplineEquicontinuity_general Suppose Assumptions 7, 8 and 9 hold, then $M_{2,N}(\theta_{N},\theta^{*})=o_{P}(N^{-\frac{1}{2}})$.
Proof.
Let $r_{N}$ be any sequence that converges to 0 at a rate slower than $s_{N}(\theta_{N},\theta^{*})$ does as $N\to\infty$. $\forall x>0,y>0$. Since
$a_{N}(r_{N},x)$ $=\frac{16Z}{\sqrt{N}}\log\mathcal{N}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)+64\mathcal{J}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)+32r_{N}\sqrt{x}+\frac{16}{\sqrt{N}}Zx$
$\leq\frac{16Z}{\sqrt{N}}\left(\frac{2\sqrt{2}L}{r_{N}}\mathcal{J}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)\right)^{2}+64\mathcal{J}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)+32r_{N}\sqrt{x}+\frac{16}{\sqrt{N}}Zx$
$\leq\left(\frac{108L^{2}Z}{\sqrt{N}r_{N}^{2}}\mathcal{J}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)+64\right)\mathcal{J}_{\infty}\left({\frac{r_{N}}{2\sqrt{2}L}},{\Theta}\right)+32r_{N}\sqrt{x}+\frac{16}{\sqrt{N}}Zx$
$\to 0,$
there exists a $N_{1}>0$, such that for every $N>N_{1}$,
$P(\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})I\left(s_{N}\left(\theta_{N},\theta^{*}\right)\leq r_{N}\right)>y)$
$\leq P\left(\sup_{\theta^{*}\in\Theta}I\left(\sigma_{\theta,N}\leq r_{N}\right)\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})\geq a_{N}(r_{N},x)\right)$
$\leq 2e^{-x},$
with the last inequality implied by Theorem 6. Similarly, $\forall x>0,y>0$, there exists a $N_{2}>0$, such that for every $N>N_{2}$, $P(-\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})I\left(s_{N}\left(\theta_{N},\theta^{*}\right)\leq r_{N}\right)>y)\leq 2e^{-x}$. Combined with Assumption 9 which implies $P(s_{N}\left(\theta_{N},\theta^{*}\right)\leq r_{N})\to 1$, we have that $\sqrt{N}M_{2,N}(\theta_{N},\theta^{*})=o_{P}(1).$
As introduced in Appendix B, van der Laan, (2023) proved the supremum-norm covering number for the class of functions $D^{(m)}([0,1]^{d})$ for $(m=1,2,\cdots)$, which implies a bound on its uniform entropy integral $J_{\infty}\left(r,D^{(m)}([0,1]^{d})\right)\asymp r^{(2m+1)/(2m+2)}(-\log r)^{d-\frac{3}{2}}$. The $m$-th order HAL-MLEs are also defined, with the uniform rate of convergence proved to be $N^{-\frac{m+1}{2m+3}}$ with up till $\log$ $N$-factors. Using these theoretical results, we provide another corollary which shows that if $\theta^{*}$ lies in $D^{(m)}([0,1]^{d})$ ($m=1,2,3,\cdots$) and $\theta_{N}$ is an estimate of $\theta^{*}$ by higher order spline HAL-MLE, then the equicontinuity result for the martingale process $M_{2,N}$ holds.
{restatable}
assumptionrepAssmpHigherOrderSpline_general Suppose that $\theta^{*}\in D^{(m)}([0,1]^{d})$ and $\theta_{N}$ is estimated by the $m$-order spline HAL-MLE $(m=1,2,\cdots)$.
{restatable}
corollaryrepCoroHighOrderSplineEquicontinuity_general Suppose Assumptions 7, 8 and 6 hold, then $M_{2,N}(\theta_{N},\theta^{*})=o_{P}(N^{-\frac{1}{2}})$.
Proof.
Let $N_{\infty}\left(r,D^{(m)}([0,1]^{d})\right)$ denote the number of spheres of size $\epsilon$ w.r.t. supremum norm that are needed to cover $D^{(m)}([0,1]^{d})$. Define $J_{\infty}\left(r,D^{(m)}([0,1]^{d})\right)$ as the corresponding supremum norm entropy integral of $D^{(m)}([0,1]^{d})$. In van der Laan, (2023), Lemma 29 shows that $J_{\infty}\left(r,D^{(m)}([0,1]^{d})\right)\asymp r^{(2m+1)/(2m+2)}(-\log r)^{d-\frac{3}{2}}$; Theorem 12 proves the uniform convergence rate $\left\|\theta_{N}-\theta^{*}\right\|_{\infty}=O_{P}(N^{-\frac{m+1}{2m+3}}\log^{c_{1}}N)$ for some $c_{1}<\infty$, which satisfies Assumption 9 that $s_{N}(\theta_{N},\theta^{*})=o_{P}(1)$. Therefore, if Assumption 7 and 8 also hold, one can apply Corollary 6 to prove that $M_{2,N}(\theta_{N},\theta^{*})=o_{P}(N^{-\frac{1}{2}})$. ∎
Appendix F Proof of Equicontinuity Result in Adaptive Designs
Recall that in Section 4
$M_{2,N}\left(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K}\right)$ $=\frac{1}{N}\sum_{i=1}^{N}\Biggl{\{}\biggl{[}D_{k}\left(\bar{Q}_{N,K}^{*}\right)\left(O_{i},C_{i}\right)-E\left[D_{k}\left(\bar{Q}_{N,K}^{*}\right)\left(O_{i},C_{i}\right)|\mathcal{H}(i)\right]\biggr{]}$
$-\biggl{[}D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)-E\left[D_{k}\left(\bar{Q}_{1,K}\right)\left(O_{i},C_{i}\right)|\mathcal{H}(i)\right]\biggr{]}\Biggr{\}}.$
where $D_{k}(\bar{Q}_{K})\left(O_{i},C_{i}\right):=\frac{g^{*}_{k,i}\left(A_{i}|C_{i}\right)}{g_{0,i}\left(A_{i}|C_{i}\right)}\left(Y_{i,K}-\bar{Q}_{K}\left(A_{i},W_{i}\right)\right)$. We apply the general results in Appendix E to prove the equicontinuity condition $M_{2,N}(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K})=o_{P}(N^{-\frac{1}{2}})$ in Theorem 3 and Corollary 3 in the context of adaptive designs.
Proof of Theorem 3 and Corollary 3.
Let $\Theta:=\bar{\mathcal{Q}}_{K}$, $f_{\theta}:=D_{k}(\bar{Q}_{K})$, $\theta^{*}:=\bar{Q}_{1,K}$, $\theta_{N}:=\bar{Q}_{N,K}$ in the general setup in Appendix E. Under Assumption 3, $\zeta<g_{0,i}<1-\zeta$ and $\zeta<g_{k,i}^{*}<1-\zeta$. Since $Y_{K}$ and $Q_{K}$ take values in the interval $[0,1]$, we have that $\sup_{\bar{Q}_{K}\in\bar{\mathcal{Q}}_{K}}\left\|D_{k}(\bar{Q}_{K})\right\|_{\infty}\leq\frac{2(1-\zeta)}{\zeta}<\infty$, $\sup_{\bar{Q}_{K,1}\in\bar{\mathcal{Q}}_{K},\bar{Q}_{K,2}\in\bar{\mathcal{Q}}_{K},\bar{Q}_{K,1}\neq\bar{Q}_{K,2}}\frac{\left\|D_{k}(\bar{Q}_{K,1})-D_{k}(\bar{Q}_{K,2})\right\|_{\infty}}{\left\|\bar{Q}_{K,1}-\bar{Q}_{K,2}\right\|_{\infty}}\leq\frac{1-\zeta}{\zeta}<\infty$. Therefore, Assumptions 7 and 8 are satisfied. Since Assumption 6 serves as an analog in our adaptive design framework to Assumption 9 in the general framework, we can utilize Corollary 6 to demonstrate that $M_{2,N}\left(\bar{Q}_{N,K}^{*},\bar{Q}_{1,K}\right)=o_{P}(N^{-\frac{1}{2}})$. Similarly, Assumption 3 parallels Assumption 6 in Corollary 6. Therefore, Corollary 3 is a specification of Corollary 6 within our adaptive design setup. ∎
Appendix G Additional Details of Simulation Setup
We estimate conditional mean functions using Super Learner (van der Laan et al.,, 2007), which incorporates Random Forest (Breiman,, 2001), XGBoost (Chen and Guestrin,, 2016) and first-order Highly Adaptive Lasso (HAL) estimator (van der Laan,, 2023). The CATE functions and their pointwise confidence intervals are estimated using the first-order HAL estimator.
Appendix H Additional Simulation Results for Section 5
In this section we present more results of our simulation study. Table 3 presents the true expected final outcome $Y_{5}$ under different adaptive designs across simulations. These values represent the “marginal” version of the context-specific parameter $\Psi_{t,k}$ discussed in Section 2.3. We denote this class of estimands as $\tilde{\Psi}_{t,k}$, with $\tilde{\Psi}_{t,RCT}$ and $\tilde{\Psi}_{t,SL}$ specifically representing the true expected final outcomes under the non-adaptive RCT and our proposed Online Superlearner design for surrogate selection, respectively.
Table 3: True expected final outcome $\tilde{\Psi}_{t,k}$ under different adaptive designs in Scenarios 1 and 2 ($t=11,21,31,41,50$).
(a) Scenario 1 $t$ $\tilde{\Psi}_{t,RCT}$ $\tilde{\Psi}_{t,1}$ $\tilde{\Psi}_{t,2}$ $\tilde{\Psi}_{t,3}$ $\tilde{\Psi}_{t,4}$ $\tilde{\Psi}_{t,5}$ $\tilde{\Psi}_{t,SL}$ 11 0 -0.010 0.073 0.124 0.143 0.131 0.047 21 0 -0.005 0.086 0.153 0.191 0.193 0.133 31 0 -0.002 0.092 0.164 0.208 0.216 0.171 41 0 0.001 0.095 0.170 0.218 0.229 0.193 50 0 0.002 0.096 0.173 0.223 0.237 0.205
(b) Scenario 2 $t$ $\tilde{\Psi}_{t,RCT}$ $\tilde{\Psi}_{t,1}$ $\tilde{\Psi}_{t,2}$ $\tilde{\Psi}_{t,3}$ $\tilde{\Psi}_{t,4}$ $\tilde{\Psi}_{t,5}$ $\tilde{\Psi}_{t,SL}$ 11 0 0.080 0.072 0.062 0.046 0.024 0.037 21 0 0.087 0.082 0.076 0.066 0.044 0.063 31 0 0.090 0.086 0.082 0.074 0.055 0.073 41 0 0.091 0.089 0.085 0.079 0.062 0.079 50 0 0.092 0.090 0.087 0.081 0.067 0.082
We also present bias, variance of TMLE point estimators and coverage probabilities of TMLE-based confidence intervals in simulations introduced in Section 5. The trajectory of regret over time is presented in Figure 4.
Table 4: Bias (×$\times 10^{-3}$) of TMLE estimators for $\Psi_{t,k}$ in Scenarios 1 and 2 ($t=11,21,31,41,50$).
(a) Scenario 1 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 -0.77 -1.84 -2.76 -2.96 -2.47 -1.15 21 -1.36 -3.34 -3.79 -3.32 -1.12 -0.01 31 -5.17 -6.20 -4.99 -3.63 -1.66 -2.36 41 -4.00 -4.19 -2.89 -1.29 -0.39 -1.31 50 -1.38 -0.06 -0.51 -0.26 0.92 -0.87
(b) Scenario 2 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 -1.56 0.26 -0.60 -0.57 -1.52 -2.17 21 -3.36 -1.87 -1.62 -2.23 -2.17 -2.83 31 -2.37 -0.51 -0.25 -0.57 -0.56 -0.83 41 -1.33 0.29 0.45 0.41 0.52 0.13 50 1.29 1.13 1.35 1.22 1.16 1.50
Table 5: Variance (×$\times 10^{-3}$) of TMLE estimators for $\Psi_{t,k}$ in Scenarios 1 and 2 ($t=11,21,31,41,50$).
(a) Scenario 1 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 3.44 7.49 5.55 4.88 4.17 3.83 21 2.45 4.21 2.93 2.36 1.85 2.23 31 1.75 3.47 2.23 1.67 1.19 1.29 41 1.28 2.66 1.74 1.31 0.89 0.90 50 1.06 2.35 1.47 1.10 0.70 0.68
(b) Scenario 2 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 3.11 4.84 4.55 4.09 3.55 3.32 21 2.23 1.51 1.49 1.45 1.50 1.94 31 1.56 0.83 0.83 0.80 0.87 1.09 41 1.18 0.60 0.59 0.59 0.67 0.76 50 0.94 0.44 0.45 0.45 0.51 0.59
Table 6: Coverage probability (%) of confidence intervals for $\Psi_{t,k}$ in Scenarios 1 and 2 ($t=11,21,31,41,50$).
(a) Scenario 1 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 96.2 95.0 95.8 96.2 96.2 95.6 21 93.8 94.0 94.8 94.6 94.2 94.2 31 94.0 92.0 94.4 95.2 95.4 94.2 41 94.4 93.2 94.6 94.4 95.2 94.0 50 95.0 93.4 96.6 96.6 94.0 95.2
(b) Scenario 2 $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 95.4 95.6 95.4 95.4 95.8 95.2 21 94.6 96.4 96.0 95.4 95.2 95.2 31 94.6 95.8 95.6 96.4 95.6 95.2 41 94.4 94.8 95.4 95.8 96.0 95.4 50 96.0 96.6 96.8 97.0 96.4 96.0
Table 7: Probability (%) of not assigning optimal individualized treatments in Scenarios 1 and 2. $p^{\textit{non-opt}}_{t}:=\frac{1}{E(t)}\sum_{i:t_{i}=t}I\left(A_{i}\neq d_{0,K}(W_{i})\right)$, where $d_{0,K}(W):=I(\bar{Q}_{0,K}(1,W)-\bar{Q}_{0,K}(0,W)>0)$
(a) Scenario 1 $t$ $p^{\textit{non-opt}}_{t,SL}$ $p^{\textit{non-opt}}_{t,1}$ $p^{\textit{non-opt}}_{t,2}$ $p^{\textit{non-opt}}_{t,3}$ $p^{\textit{non-opt}}_{t,4}$ $p^{\textit{non-opt}}_{t,5}$ $p^{\textit{non-opt}}_{t,RCT}$ 11 28.4 51.4 39.7 30.0 20.1 15.9 49.9 21 20.2 50.8 40.1 30.0 20.6 14.7 49.8 31 17.5 50.1 39.8 29.8 20.3 14.1 50.0 41 14.9 49.9 40.0 30.1 19.7 13.0 50.1 50 14.9 50.1 39.8 29.6 19.9 13.2 50.5
(b) Scenario 2 $t$ $p^{\textit{non-opt}}_{t,SL}$ $p^{\textit{non-opt}}_{t,1}$ $p^{\textit{non-opt}}_{t,2}$ $p^{\textit{non-opt}}_{t,3}$ $p^{\textit{non-opt}}_{t,4}$ $p^{\textit{non-opt}}_{t,5}$ $p^{\textit{non-opt}}_{t,RCT}$ 11 15.3 14.0 14.6 16.2 19.5 30.9 50.3 21 13.4 12.7 13.1 14.4 16.6 24.5 50.2 31 12.8 12.3 12.7 13.9 15.7 21.6 49.4 41 11.9 11.7 11.9 13.0 15.0 19.3 50.8 50 12.1 11.5 12.0 12.6 14.4 18.6 50.5
Table 8: Average regret of not selecting optimal individualized treatments in Scenarios 1 and 2.
(a) Scenario 1 $t$ $R_{t,SL}$ $R_{t,1}$ $R_{t,2}$ $R_{t,3}$ $R_{t,4}$ $R_{t,5}$ $R_{t,RCT}$ 11 0.154 0.348 0.240 0.160 0.099 0.083 0.344 21 0.102 0.340 0.241 0.159 0.101 0.080 0.341 31 0.088 0.337 0.241 0.158 0.100 0.078 0.342 41 0.078 0.331 0.238 0.156 0.093 0.071 0.343 50 0.077 0.336 0.240 0.155 0.096 0.074 0.348
(b) Scenario 2 $t$ $R_{t,SL}$ $R_{t,1}$ $R_{t,2}$ $R_{t,3}$ $R_{t,4}$ $R_{t,5}$ $R_{t,RCT}$ 11 0.030 0.027 0.028 0.029 0.035 0.064 0.121 21 0.026 0.026 0.026 0.027 0.030 0.048 0.121 31 0.025 0.025 0.026 0.026 0.028 0.040 0.119 41 0.024 0.025 0.025 0.025 0.027 0.035 0.122 50 0.023 0.024 0.024 0.025 0.027 0.033 0.121
Figure 4: Regret at time $t$ for different design strategies in Scenario 1 (left) and Scenario 2 (right). This figure includes seven curves: five curves correspond to regrets from adaptive designs using outcomes from $Y_{1}$ to $Y_{5}$, respectively; one curve illustrates the regret for an adaptive design employing an Online Superlearner to evaluate and utilize surrogates; and the last curve represents the regret of a fixed design, which assigns equal probabilities to each treatment group.
Appendix I Additional Simulation Results for Section 6
This section presents more results for ADAPT-R simulation in Section 6. Figure 5 shows the estimated CATE functions of Navigator v.s. SMS+CCT to generate data in ADAPT-R simulation. Table 9 and 10 show bias, variance of TMLE point estimators and coverage probabilities of confidence intervals. Table 11 presents TMLE’s performance of estimating ATE on final outcome and mean final outcome of optimal treatment rules based on surrogate outcomes and final outcomes after the experiment ends. Table 12 shows the true expected final outcome $\tilde{\Psi}_{t,k}$ in ADAPT-R simulation.
Figure 5: Estimated Conditional Average Treatment Effect (CATE) functions of Navigator v.s. SMS+CCT in Phase 2, plotted against initial retention strategy (grids) and time to lapse after initial strategy in Phase 1 (x-axis). Each curve corresponds to the CATE function of a candidate surrogate, ranging from $Y_{1}$ to $Y_{5}$.
Table 9: Bias and Variance of TMLE estimators for $\Psi_{t,k}$ in ADAPT-R trial simulations.
(a) Bias (×$\times 10^{-3}$) $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 -0.33 -0.80 -0.06 -0.17 -0.40 -0.27 13 0.06 -0.69 -0.25 0.07 -0.60 -0.18 15 0.10 -0.11 0.25 0.35 -0.11 0.13 17 0.10 -0.03 0.12 0.33 -0.17 -0.01 19 -0.14 -0.21 -0.12 -0.02 -0.28 -0.25
(b) Variance (×$\times 10^{-3}$) $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 0.17 0.25 0.23 0.21 0.20 0.19 13 0.15 0.18 0.21 0.20 0.20 0.19 15 0.12 0.13 0.14 0.16 0.16 0.15 17 0.10 0.11 0.12 0.14 0.13 0.13 19 0.10 0.10 0.11 0.13 0.12 0.12
Table 10: True target estimand $\Psi_{t,k}$ had participants enrolled from $1,\cdots,t-k$ were adaptively treated by a candidate surrogate $Y_{k}$ in ADAPT-R trial simulations ($t=11,13,15,17,19$).
(a) $\Psi_{t,k}$ $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 0.538 0.541 0.540 0.538 0.538 0.538 13 0.547 0.556 0.552 0.551 0.549 0.547 15 0.557 0.570 0.567 0.564 0.562 0.560 17 0.562 0.577 0.575 0.573 0.570 0.568 19 0.564 0.579 0.576 0.574 0.572 0.570
(b) Coverage probability (%) $t$ $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ 11 95.8 95.9 96.2 96.7 95.8 95.4 13 93.6 94.8 92.4 94.7 94.1 93.4 15 94.2 95.0 93.8 94.9 94.7 93.7 17 94.0 94.8 95.1 95.3 94.7 94.8 19 93.6 95.5 95.5 94.9 94.9 93.7
Table 11: Performance of TMLE estimators for other causal estimands after the experiments in ADAPT-R trial simulation.
Estimand Notation Truth Bias (×$\times 10^{-3}$) Var (×$\times 10^{-3}$) Coverage (%) Average Treatment Effect on Final Outcome $Y_{5}$ $\psi_{n,5,T}^{\text{ATE}}$0.015 1.14 0.34 94.8 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{1}$ $\psi_{d^{*}_{n,1,T}}$ 0.590 0.18 0.13 96.2 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{2}$ $\psi_{d^{*}_{n,2,T}}$ 0.592 0.44 0.15 93.4 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{3}$ $\psi_{d^{*}_{n,3,T}}$ 0.593 0.44 0.17 91.8 Expected Final Outcome $Y_{5}$ under OTR of Surrogate $Y_{4}$ $\psi_{d^{*}_{n,4,T}}$ 0.595 0.04 0.16 92.0 Expected Final Outcome $Y_{5}$ under OTR of $Y_{5}$ $\psi_{d^{*}_{n,5,T}}$ 0.593 1.03 0.21 94.8
Table 12: True expected final outcome $\tilde{\Psi}_{t,k}$ under different designs in ADAPT-R trial simulation.
t $\Psi_{t,RCT}$ $\Psi_{t,1}$ $\Psi_{t,2}$ $\Psi_{t,3}$ $\Psi_{t,4}$ $\Psi_{t,5}$ $\Psi_{t,SL}$
11 0.538 0.541 0.541 0.538 0.538 0.538 0.538
13 0.547 0.553 0.550 0.550 0.549 0.547 0.551
15 0.557 0.567 0.564 0.562 0.561 0.560 0.563
17 0.562 0.575 0.573 0.571 0.570 0.567 0.571
19 0.564 0.576 0.574 0.573 0.572 0.570 0.572
Appendix J An Extension of Our Proposed CARA Design
In the adaptive design with the Online Superlearner, as investigated in previous sections, a conservative treatment randomization strategy is employed, ensuring that adaptations based on surrogate outcomes, as selected by the Online Superlearner, are initiated only subsequent to the observation of the primary outcome at $t=K+1$. However, there may be scenarios where experimenters wish to adapt treatment randomization probabilities at an earlier stage. To address this, we propose a modified strategy in which a plausible surrogate outcome (e.g., $Y_{k}$) is designated as a temporary primary outcome. This surrogate serves as the foundation for implementing our adaptive design that treats $Y_{k}$ as the primary outcome and enabling the adaptation of treatment randomization probabilities as early as $t=k+1$. As the experiment progresses, subsequent surrogate outcomes can be sequentially established as new temporary primary outcomes until the true primary outcome is observed. If one has prior knowledge or external datasets to infer the relationship between surrogates and primary outcomes, one can apply other methods to start adapting treatment randomization probabilities earlier. However, that usually depends on generalization assumptions or pre-specified models between surrogates and primary outcomes. We note that whatever strategy the experimenter chooses, our framework with Online Superlearner provides a tool for experimenters to evaluate how outcomes would be different under other designs that could have been implemented, and dynamically utilizes the most useful design within the ongoing sequential decision-making environment.